### v2

# From Matching to Generation: A Survey on Generative Information Retrieval

Abstract.
Information Retrieval (IR) systems are crucial tools for users to access information, which have long been dominated by traditional methods relying on similarity matching. With the advancement of pre-trained language models, generative information retrieval (GenIR) emerges as a novel paradigm, attracting increasing attention. Currently, research in GenIR can be categorized into two aspects: generative document retrieval (GR) and reliable response generation. GR leverages the generative model’s parameters for memorizing documents, enabling retrieval by directly generating relevant document identifiers without explicit indexing. Reliable response generation, on the other hand, employs language models to directly generate the information users seek, breaking the limitations of traditional IR in terms of document granularity and relevance matching, offering more flexibility, efficiency, and creativity, thus better meeting practical needs. This paper aims to systematically review the latest research progress in GenIR. We will summarize the advancements in GR regarding model training and structure, document identifier, incremental learning, etc., as well as progress in reliable response generation in aspects of internal knowledge memorization, external knowledge augmentation, etc. We also review the evaluation, challenges and future developments in GenIR systems. This review aims to offer a comprehensive reference for researchers, encouraging further development in the GenIR field.  [^1] Github Repository: https://github.com/RUC-NLPIR/GenIR-Survey
Generative Information Retrieval; Generative Document Retrieval; Reliable Response Generation
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: TOISccs: Information systems Retrieval models and ranking
1. Introduction
Figure 1. Exploring IR Evolution: From Traditional to Generative Methods - This diagram illustrates the shift from traditional similarity-based document matching (a) to GenIR techniques. Current GenIR methods can be categorized into two types: generative retrieval (b), which retrieves documents by directly generating relevant DocIDs constrained by a DocID prefix tree; and response generation (c), which directly generates reliable and user-centric answers.
In today’s digital landscape, information retrieval (IR) systems are crucial for navigating the vast sea of online information. From using search engines such as Google (Google, 2023a), Bing (Microsoft, 2023a), and Baidu (OpenAI, 2022), to engaging with question-answering or dialogue systems like ChatGPT (OpenAI, 2022) and Bing Chat (Microsoft, 2023b), and discovering content via recommendation platforms like Amazon (Amazon, 2023) and YouTube (Google, 2023b), IR technologies are integral to our everyday online experiences. These systems are not only reliable but also play a key role in spreading knowledge and ideas globally.
Traditional IR systems primarily rely on sparse retrieval methods based on word-level matching. These methods, which include Boolean Retrieval (Salton et al., 1983), BM25 (Robertson and Zaragoza, 2009), SPLADE (Formal et al., 2021), and UniCOIL (Lin and Ma, 2021), establish connections between vocabulary and documents, offering high retrieval efficiency and robust system performance. With the rise of deep learning, dense retrieval methods such as DPR (Karpukhin et al., 2020) and ANCE (Xiong et al., 2020), based on the bidirectional encoding representations from the BERT model (Kenton and Toutanova, 2019), capture the deep semantic information of documents, significantly improving retrieval precision. Although these methods have achieved leaps in accuracy, they rely on large-scale document indices (Douze et al., 2024; Malkov and Yashunin, 2020) and cannot be optimized in an end-to-end way. Moreover, when people search for information, what they really need is a precise and reliable answer. This document ranking list-based IR approach still requires users to spend time summarizing their required answers, which is not ideal enough for information seeking (Metzler et al., 2021).
With the development of Transformer-based pre-trained language models such as T5 (Raffel et al., 2020), BART (Lewis et al., 2020a), and GPT (Radford et al., 2018), they have demonstrated their strong text generation capabilities. In recent years, large language models (LLMs) have brought about revolutionary changes in the field of AI-generated content (AIGC) (Zhao et al., 2023; Cao et al., 2023). Based on large pre-training corpora and advanced training techniques like RLHF (Christiano et al., 2017), LLMs (OpenAI, 2022; Touvron et al., 2023b; Bai et al., 2023; Jiang et al., 2023b) have made significant progress in natural language tasks, such as dialogue (OpenAI, 2022; Thoppilan et al., 2022) and question answering (Liu et al., 2023a; Qin et al., 2023a). The rapid development of LLMs is transforming IR systems, giving rise to a new paradigm of generative information retrieval (GenIR), which achieves IR goals through generative approaches.
As envisioned by Metzler et al. (Metzler et al., 2021), in order to build an IR system that can respond like a domain expert, the system should not only provide accurate responses but also include source citations to improve credibility and transparency of the results. To achieve this, GenIR models must possess both sufficient memorized knowledge and the ability to recall the associations between knowledge and corresponding documents. Current research in GenIR is primarily focused on two main patterns: (1) generative document retrieval (GR), which involves retrieving documents by generating their identifiers; and (2) reliable response generation, which entails directly generating user-centric responses through strategies that enhance their reliability. These two patterns are the central topics of this survey.
Generative document retrieval, a new retrieval paradigm based on generative models, is garnering increasing attention. This approach leverages the parametric memory of generative models to directly generate document identifiers (DocIDs) related to the documents (Cao et al., 2021b; Tay et al., 2022; Wang et al., 2022a; Zhou et al., 2022). Figure 1 illustrates this transition, where traditional IR systems match queries to documents based on indexed database (Figure 1(a)), while generative methods use language models to retrieve by directly generating relevant document identifiers (Figure 1(b)). Specifically, GR assigns a unique identifier to each document, which can be numeric-based or text-based, and then trains a generative retrieval model to learn the mapping from queries to the relevant document DocIDs. This allows the model to index documents using its internal parameters. During inference, GR models use constrained beam search to limit the generated DocIDs to be valid within the corpus, ranking them based on generation probability to produce a ranked list of DocIDs. This eliminates the need for large-scale document indexes in traditional methods, enabling end-to-end training of the model.
Recent studies on generative retrieval have delved into model training and structure (Tay et al., 2022; Wang et al., 2022a; Zhou et al., 2023c, a), document identifier design (Cao et al., 2021b; Tay et al., 2022; Sun et al., 2023d; Yang et al., 2023c), continual learning on dynamic corpora (Mehta et al., 2022; Kishore et al., 2023; Guo et al., 2024), downstream task adaptation (Chen et al., 2022a, b; Li et al., 2024a), multi-modal generative retrieval (Zhang et al., 2023g; Long et al., 2024; Li et al., 2024e), and generative recommender systems (Geng et al., 2022; Wang et al., 2023e; Rajput et al., 2023). The progress in GR is shifting retrieval systems from matching to generation. It has also led to the emergence of workshops (Bénédict et al., 2023) and tutorials (Tang et al., 2023a). However, there is currently no comprehensive review that systematically organizes the research, challenges, and prospects of this emerging field.
Reliable response generation is also a promising direction in the IR field, offering user-centric and accurate answers that directly meet users’ needs. Since LLMs are particularly adept at following instructions (Zhao et al., 2023), capable of generating customized responses, and can even cite the knowledge sources (Nakano et al., 2021; Qian et al., 2023b), making direct response generation a new and intuitive way to access information (Sallam et al., 2023; Gienapp et al., 2023; White, 2023). As illustrated in Figure 1, the generative approach marks a significant shift from traditional IR systems, which return a ranked list of documents (as shown in Figure 1(a,b)). Instead, response generation methods (depicted in Figure 1(c)) offer a more dynamic form of information access by directly generating detailed, user-centric responses, thereby providing a richer and more immediate understanding of the information need behind the users’ queries.
However, the responses generated by language models may not always be reliable. They have the potential to generate irrelevant answers (Hersh, 2023), contradict factual information (Ji et al., 2023a; Huang et al., 2023d), provide outdated data (Vu et al., 2023), or generate toxic content (Huang et al., 2023e; Sun et al., 2024). Consequently, these limitations render them unsuitable for many scenarios that require accurate and up-to-date information. To address these challenges, the academic community has developed strategies across four key aspects: enhancing internal knowledge (Brown et al., 2020; Scao et al., 2022; Touvron et al., 2023a; Lee et al., 2022a; Dong et al., 2022; Chuang et al., 2023; Sun et al., 2020; Ke et al., 2023; Meng et al., 2022); augmenting external knowledge (Lewis et al., 2020b; Wang et al., 2024c; Asai et al., 2023; Yao et al., 2022; Nakano et al., 2021; Jiang et al., 2023f, e; Schick et al., 2023); generating responses with citation (Weller et al., 2023; Lee et al., 2023d; Nakano et al., 2021; Li et al., 2023j, 2024d); and improving personal information assistance (Liu et al., 2020; Wang et al., 2023c; Xu et al., 2023c; Li and Tuzhilin, 2019). Despite these efforts, there is still a lack of a systematic review that organizes the existing research under this new paradigm of generative information access.
{forest} forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=draw-leaf, rounded corners, align=left, text centered, minimum width=5em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=1.2pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=0text width=7.5em,fill=fill_0,draw=draw_0,font=,, where level=1text width=10em,fill=fill_1,draw=draw_1,font=,, where level=2text width=10.4em,fill=fill_2,draw=draw_2,font=,, where level=3font=,, [ GenIR System, minimum height=2.8em[ Generative Document
Retrieval (Sec 3)[ Model Training and
Structure (Sec 3.1)[ Training (3.1.1): DSI (Tay et al., 2022), DynamicRetriever (Zhou et al., 2023b), NCI (Wang et al., 2022a), DSI-QG (Zhuang et al., 2022),
Chen et al. (Chen et al., 2023a), LTRGR (Li et al., 2023d), GenRRL (Zhou et al., 2023a), DGR (Li et al., 2024f), ListGR (Tang et al., 2024b),, leaf, text width=31.8em ] [ Structure (3.1.2): NCI (Wang et al., 2022a), TOME (Ren et al., 2023), NP Decoding (Lee et al., 2023b), MEVI (Zhang et al., 2023e),
DiffusionRet (Qiao et al., 2023), GDR (Yuan et al., 2024), Self-Retrieval (Tang et al., 2024a), PAG (Zeng et al., 2024), leaf, text width=31.8em ] ] [ Document Identifier
(Sec 3.2)[ Numeric (3.2.1): DSI (Tay et al., 2022), DynamicRetriever (Zhou et al., 2023b), Ultron (Zhou et al., 2022), GenRet (Sun et al., 2023d),
Tied-Atomic (Nguyen and Yates, 2023), MEVI (Zhang et al., 2023e), LMIndexer (Jin et al., 2023), ASI (Yang et al., 2023c), RIPOR (Zeng et al., 2023), leaf, text width=31.8em ] [ Text (3.2.2): GENRE (Cao et al., 2021b), SEAL (Bevilacqua et al., 2022), Ultron (Zhou et al., 2022), LLM-URL (Ziems et al., 2023),
UGR (Chen et al., 2023e), MINDER (Li et al., 2023e), AutoTSG (Zhang et al., 2023b), SE-DSI (Zhang et al., 2023b), NOVO (Wang et al., 2023l), GLEN (Lee et al., 2023a), leaf, text width=31.8em ] ] [ Incremental Learning
(Sec 3.3)[ DSI++ (Mehta et al., 2022), IncDSI (Kishore et al., 2023), CLEVER (Chen et al., 2023d), CorpusBrain++ (Guo et al., 2024), leaf, text width=31.8em ] ] [ Downstream Task
Adaptation (Sec 3.4)[ Separate Training (3.4.1): GERE (Chen et al., 2022a), CorpusBrain (Chen et al., 2022b), GMR (Lee et al., 2022c), DearDR (Thorne, 2022),
CodeDSI (Nadeem et al., 2022), UGR (Chen et al., 2023e), GCoQA (Li et al., 2023c), Re3val (Song et al., 2024), leaf, text width=31.8em ] [ Joint Training (3.4.2): UniGen (Li et al., 2024g), CorpusLM (Li et al., 2024a), leaf, text width=31.8em ] [ Multi-Modal (3.4.3): IRGen (Zhang et al., 2023g), GeMKR (Long et al., 2024), GRACE (Li et al., 2024e), leaf, text width=31.8em ] ] [ Generative Recommender
Systems (Sec 3.5), text width=11.9em[ P5 (Geng et al., 2022), TIGER (Rajput et al., 2023), SEATER (Si et al., 2023), IDGenRec (Tan et al., 2024b), LC-Rec (Zheng et al., 2023),
ColaRec (Wang et al., 2024b), leaf, text width=30.3em ] ] ] [ Reliable Response
Generation (Sec 4) [ Internal Knowledge
Memorization (Sec 4.1) [ Structure (4.1.1): Model Scaling: GPT3 (Brown et al., 2020), BLOOM (Scao et al., 2022), LLaMA (Touvron et al., 2023a),
Model Structure: PaLM (Chowdhery et al., 2023), Mixtral 8x7B (Jiang et al., 2024), leaf, text width=31.8em ] [ Training and Inference (4.1.2): Training: Sadeq et al. (Sadeq et al., 2023) FactTune (Ovadia et al., 2023a);
Inference: GenRead (Yu et al., 2022), RECITE (Sun et al., 2022b), DoLa (Chuang et al., 2023), leaf, text width=31.8em] [ Knowledge Updating (4.1.3): Incremental Learning: Ernie 2.0 (Sun et al., 2020), DAP (Ke et al., 2023);
DynaInst (Mok et al., 2023), Knowledge Editing: KE (Cao et al., 2021a), MEND (Mitchell et al., 2022), ROME (Meng et al., 2022) , leaf, text width=31.8em ] ] [ External Knowledge
Augmentation (Sec 4.2) [ Retrieval (4.2.1): Sequential: RAG (Lewis et al., 2020b), RRR (Ma et al., 2023a), ARL2 (Zhang et al., 2024b);
Branching: TOC (Kim et al., 2023), BlendFilter (Wang et al., 2024d), REPLUG (Shi et al., 2023b);
Conditional: SKR (Wang et al., 2023d), Self-DC (Wang et al., 2024c), Rowen (Ding et al., 2024);
Loop: Iter-RetGen (Shao et al., 2023), IR-COT (Trivedi et al., 2022), FLARE (Jiang et al., 2023e), Self-RAG (Asai et al., 2023) , leaf, text width=31.8em ] [ Tool (4.2.2): Search Engine: ReAct (Yao et al., 2022), WebGPT (Nakano et al., 2021);
Knowledge Graph: StructGPT (Jiang et al., 2023f), ToG (Sun et al., 2023c), RoG (Luo et al., 2024);
API-based Tools: Toolformer (Schick et al., 2023), ToolLLM (Qin et al., 2023b), AssistGPT (Gao et al., 2023b);
Model-based Tools: HuggingGPT (Shen et al., 2023), Visual ChatGPT (Wu et al., 2023b) , leaf, text width=31.8em ] ] [ Generating Response
with Citation (Sec 4.3) [ Direct Citation (4.3.1): According-to Prompting (Weller et al., 2023), IFL (Lee et al., 2023d), Fierro et al. (Fierro et al., 2024),
Credible without Credit (Peskoff and Stewart, 2023), 1-PAGER (Jain et al., 2023), Khalifa et al. (Khalifa et al., 2024) , leaf, text width=31.8em ] [ Retrieval-based Citation (4.3.2): WebGPT (Nakano et al., 2021), WebBrain (Qian et al., 2023b), RARR (Gao et al., 2023a),
SearChain (Xu et al., 2023a), LLatrieval (Gao et al., 2023c), VTG (Sun et al., 2023a), CEG (Li et al., 2024b), APO (Li et al., 2024d) , leaf, text width=31.8em ] ] [ Personal Information
Assistant (Sec 4.4) [ Personalized Dialogue (4.4.1): Zhang et al. (Zhang et al., 2018), $\mathcal{P}^{2}$ot (Liu et al., 2020), Wu et al. (Wu et al., 2021),
SAFARI (Wang et al., 2023c), Personalized Soups (Jang et al., 2023a), OPPU (Tan et al., 2024c) , leaf, text width=31.8em ] [ Domain-specific (4.4.2): Healthcare: Zhongjing (Liu et al., 2023e), Mental-LLM (Xu et al., 2023c);
Academic: RevGAN (Li and Tuzhilin, 2019), Pearl (Mysore et al., 2023); Education: EduChat (Dan et al., 2023); Recipe, Robot, etc. , leaf, text width=31.8em ] ] ] [ Evaluation (Sec 5)[ Generative Document
Retrieval (Sec 5.1)[ Metrics (5.1.1): Recall, MRR (Craswell, 2009), R-Precision, MAP, nDCG (Järvelin and Kekäläinen, 2002), leaf, text width=31.8em ] [ Benchmarks (5.1.2): MS MARCO (Nguyen et al., 2016), NQ (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), KILT (Petroni et al., 2021),
TREC DL 19 & 20 (Craswell et al., 2020, 2021), DynamicIR (Yoon et al., 2023), Liu et al. (Liu et al., 2023g), ExcluIR (Zhang et al., 2024c), leaf, text width=31.8em ] [ Analysis (5.1.3): Chen et al. (Chen et al., 2023a), Pradeep et al. (Pradeep et al., 2023), Liu et al. (Liu et al., 2023g),
Wu et al. (Wu et al., 2024b), leaf, text width=31.8em ] ] [ Reliable Response
Generation (Sec 5.2)[ Metrics (5.2.1): Rule-based: EM, BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), Perplexity;
Model-based: BERTScore (Zhang et al., 2019), BLEURT (Sellam et al., 2020), GPTScore (Fu et al., 2023), FActScore (Min et al., 2023);
Human Evaluation: Comprehensibility, Relevance, Fluency , leaf, text width=31.8em ] [ Benchmarks (5.2.2): General: MMLU (Hendrycks et al., 2021), BIG-bench (Srivastava et al., 2022), LLM-Eval (Lin and Chen, 2023);
Tool: API-Bank (Li et al., 2023i), ToolBench (Qin et al., 2023b);
Factuality: TruthfulQA (Lin et al., 2022), ALCE (Gao et al., 2023c), HaluEval (Li et al., 2023a);
Real Time: RealTime QA (Kasai et al., 2023), FreshQA (Vu et al., 2023);
Trustworthy: SafetyBench (Zhang et al., 2023a), TrustGPT (Huang et al., 2023e), TrustLLM (Sun et al., 2024) , leaf, text width=31.8em ] ] ] [ Challenges and
Prospects (Sec 6)[ Generative Document
Retrieval (Sec 6.1)[ Scalability (6.1.1); Dynamic Corpora (6.1.2); Document Representation (6.1.3);
Efficiency (6.1.4) , leaf, text width=31.8em ] ] [ Reliable Response
Generation (Sec 6.2)[ Accuracy and Factuality (6.2.1); Real-time Property (6.2.2);
Bias and Fairness (6.2.3); Privacy and Security (6.2.4) , leaf, text width=31.8em ] ] [ Unified Framework
(Sec 6.3)[ Unified Framework for Retrieval and Generation (6.3.1);
Towards End2end Framework for Various IR Tasks (6.3.2) , leaf, text width=31.8em ] ] ] ]
Figure 2. Taxonomy of research on generative information retrieval: investigating generative document retrieval, reliable response generation, evaluation, challenges and prospects.
This review will systematically review the latest research progress and future developments in the field of GenIR, as shown in Figure 2, which displays the classification of research related to the GenIR system. We will introduce background knowledge in Section 2, generative document retrieval technologies in Section 3, direct information accessing with generative language models in Section 4, evaluation in Section 5, current challenges and future directions in Section 6, respectively. Section 7 will summarize the content of this review. This article is the first to systematically organize the research, evaluation, challenges and prospects of generative IR, while also looking forward to the potential and importance of GenIR’s future development. Through this review, readers will gain a deep understanding of the latest progress in developing GenIR systems and how it shapes the future of information access.
The main contribution of this survey is summarized as follows:
  • First comprehensive survey on generative information retrieval (GenIR): This survey is the first to comprehensively organize the techniques, evaluation, challenges, and prospects on the emerging field of GenIR, providing a deep understanding of the latest progress in developing GenIR systems and its future in shaping information access.
  • Systematic categorization and in-depth analysis: The survey offers a systematic categorization of research related to GenIR systems, including generative document retrieval, reliable response generation. It provides an in-depth analysis of each category, covering model training and structure, document identifier, etc. in generative document retrieval; internal knowledge memorization, external knowledge enhancement, etc. for reliable response generation.
  • Comprehensive review of evaluation metrics and benchmarks: The survey reviews a range of widely used evaluation metrics and benchmark datasets for accessing GenIR methods, alongside analysis on the effectiveness and weaknesses of existing GenIR methods.
  • Discussions of current challenges and future directions: The survey identifies and discusses the current challenges faced in the GenIR field. We also provide potential solutions for each challenge and outline future research directions for building GenIR systems.
2. Background and Preliminaries
Information retrieval techniques aim at efficiently obtaining, processing, and understanding information from massive data. Technological advancements have driven the evolution of retrieval methods, from traditional keyword-based sparse retrieval, to deep learning-based dense retrieval, and further to recently emerged generative retrieval and large language models. Each advancement enhances retrieval accuracy and efficiency, catering to the complex and diverse query needs of users.
2.1. Traditional Information Retrieval
Sparse Retrieval. In the field of traditional information retrieval, sparse retrieval techniques implement fast and accurate document retrieval through the inverted index method. Inverted indexing technology maps each unique term to a list of all documents containing that term, providing an efficient means for information retrieval in large document collections. Among these methods, TF-IDF (Term Frequency-Inverse Document Frequency) (Ramos et al., 2003) is a particularly important statistical tool used to assess the importance of a word in a document collection, thereby widely applied in various traditional retrieval systems.
The core of sparse retrieval technology lies in evaluating the relevance between documents and user queries. Specifically, given a document collection $\mathcal{D}$ and a user query $q$, traditional information retrieval systems identify and retrieve information by calculating the relevance $\mathcal{R}$ between document $d$ and query $q$. This relevance evaluation typically relies on the similarity measure between document $d$ and query $q$, as shown below:
(1) $\mathcal{R}(q,d)=\sum_{t\in q\cap d}\text{tf-idf}(t,d)\cdot\text{tf-idf}(t,q),$
[…]
Dense Retrieval. The advent of pre-trained language models like BERT (Kenton and Toutanova, 2019) has revolutionized information retrieval, leading to the development of dense retrieval methods, like DPR (Karpukhin et al., 2020), ANCE (Xiong et al., 2020), E5 (Wang et al., 2022b), SimLM (Wang et al., 2023h). Unlike traditional sparse retrieval, these methods leverage Transformer-based encoders to create dense vector representations for both queries and documents. This approach enhances the capability to grasp the underlying semantics, thereby improving retrieval accuracy.
The core of dense retrieval lies in converting documents and queries into vector representations. Given document $d$ and query $q$ and their vector representations $\mathbf{v}_{q}$, each document $d$ is transformed into a dense vector $\mathbf{v}_{d}$ through a pre-trained language model, similarly, query $q$ is transformed into vector $\mathbf{v}_{q}$. Specifically, we can use encoder functions $E_{d}(\cdot)$ and $E_{q}(\cdot)$ to represent the encoding process for documents and queries, respectively:
(2) $\mathbf{v}_{d}=E_{d}(d),\quad\mathbf{v}_{q}=E_{q}(q),$
where $E_{d}(\cdot)$ and $E_{q}(\cdot)$ can be the same model or different models optimized for specific tasks.
Dense retrieval methods evaluate relevance by calculating the similarity between the query vector and document vector, which can be measured by cosine similarity, expressed as follows:
(3) $\mathcal{R}(q,d)=\text{cos}(\mathbf{v}_{q},\mathbf{v}_{d})=\frac{\mathbf{v}_{q}\cdot\mathbf{v}_{d}}{|\mathbf{v}_{q}||\mathbf{v}_{d}|},$
where $\mathbf{v}_{q}\cdot\mathbf{v}_{d}$ represents the dot product of query vector $\mathbf{v}_{q}$ and document vector $\mathbf{v}_{d}$, and $|\mathbf{v}_{q}|$ and $|\mathbf{v}_{d}|$ respectively represent the magnitudes of the query and document vector. Finally, documents are ranked based on these similarity scores to identify the most relevant ones for the user.
2.2. Generative Retrieval
With significant progress of language models, generative retrieval (Metzler et al., 2021; Tay et al., 2022) has emerged as a new direction in the field of information retrieval. Unlike traditional index-based retrieval methods, generative retrieval relies on pre-trained generative models, such as the T5 (Raffel et al., 2020) and BART (Lewis et al., 2020a), to directly generate document identifiers (DocIDs) related to the query, thereby achieving end-to-end retrieval without relying on large-scale pre-built document indices.
Formally, given a query $q$, a generative retrieval model aims to directly generate the document identifier $d^{\prime}$ related to $q$. The model can be represented as a conditional probability distribution, where the probability of generating document identifier $d^{\prime}$ given query $q$ can be expressed as:
(4) $\mathcal{R}(q,d)=P(d^{\prime}|q;\theta)=\prod_{i=1}^{T}P(d^{\prime}_{i}|d^{\prime}_{<i},q;\theta),$
where $T$ is the length of the generated document identifier $d^{\prime}$, $d^{\prime}_{i}$ represents the token at position $i$, $d^{\prime}_{<i}$ represents the sequence of tokens generated before position $i$, and $\theta$ represents the model parameters.
Generative retrieval models demonstrate flexibility and deep semantic understanding capabilities. Compared to traditional methods, generative retrieval achieves end-to-end optimization of the retrieval process by directly generating DocIDs. Additionally, it reduces reliance on external indexing, lowering the system’s demand for storage resources. However, its performance and efficiency are limited by model capacity and computational resources, indicating that future research might focus on finding more efficient model architectures and training methods to better balance performance, resource consumption, and response speed.
2.3. Large Language Models
The evolution of Large Language Models (LLMs) marks a significant leap in natural language processing (NLP), rooted from early statistical and neural network-based language models (Zhu et al., 2023). These models, through pre-training on vast text corpora, learned deep semantic features of language, greatly enriching the understanding of text. Subsequently, generative language models, most notably the GPT series (Radford et al., 2018, 2019; Brown et al., 2020), significantly improved text generation and understanding capabilities with improved model size and number of parameters.
LLMs can be mainly divided into two categories: encoder-decoder models and decoder-only models. Encoder-decoder models, like T5 (Raffel et al., 2020) and BART (Lewis et al., 2020a), convert input text into vector representations through their encoder, then the decoder generates output text based on these representations. This model perspective treats various NLP tasks as text-to-text conversion problems, solving them through text generation. On the other hand, decoder-only models, like the GPT (Radford et al., 2018) and GPT-2 (Radford et al., 2019), rely entirely on the Transformer decoder, generating text step by step through the self-attention mechanism. The introduction of GPT-3 (Brown et al., 2020), with its 175 billion parameters, marked a significant milestone in this field and led to the creation of models like InstructGPT (Ouyang et al., 2022), Falcon (Penedo et al., 2023), PaLM (Chowdhery et al., 2023) and LLaMA (Touvron et al., 2023a). These models, all using a decoder-only architecture, trained on large-scale datasets, have shown astonishing language processing capabilities (Zhao et al., 2023).
In the field of information retrieval, LLMs utilize two main strategies: In-Context Learning (ICL) and parameter-efficient fine-tuning (Zhu et al., 2023). ICL leverages the knowledge pre-trained in the model to understand and handle new tasks, requiring only a series of relevant examples and task descriptions to be provided to the model, enabling it to address specific information retrieval tasks without additional training (Brown et al., 2020). This method’s flexibility and cost-effectiveness are extremely high, allowing the model to quickly adapt to various scenarios. On the other hand, parameter-efficient fine-tuning, such as LoRA (Hu et al., 2022), PEFT (Ding et al., 2023) and QLoRA (Dettmers et al., 2023) technologies, adjusts a very small part of the model’s parameters to adapt to specific tasks, avoiding the need to retrain the entire model. These techniques, by fine-tuning key parts of the model, significantly reduce reliance on computational resources while maintaining performance.
Incorporating these strategies, LLMs have demonstrated exceptional potential across various information retrieval tasks, including but not limited to query rewriting, document retrieval, result re-ranking, document reading comprehension, and agent-based search mechanisms (Zhu et al., 2023, 2024). The application of LLMs extends beyond enhancing the relevance of search outcomes; they are instrumental in directly generating the precise information sought by users. This capability marks a significant step towards a new era of generative information retrieval. In this era, the retrieval process is not solely about locating existing information but also about creating new content that meets the specific needs of users. This feature is especially advantageous in situations where users might not know how to phrase their queries or when they are in search of complex and highly personalized information, scenarios where traditional matching-based methods fall short.
3. Generative Document Retrieval: From Similarity Matching to Generating Document Identifiers
Figure 3. Timeline of research in generative retrieval: focus on model training and structure, document identifier design, incremental learning and downstream task adaptation.
In recent advancements in AIGC, generative retrieval (GR) has emerged as an promising approach in the field of information retrieval, garnering increasing interest from the academic community. Figure 3 showcases a timeline of the GR methods. Initially, GENRE (Cao et al., 2021b) proposed to retrieve entities by generating their unique names through constrained beam search via a pre-built entity prefix tree, achieving advanced entity retrieval performance. Subsequently, Metzler et al. (Metzler et al., 2021) envisioned a model-based information retrieval framework aiming to combine the strengths of traditional document retrieval systems and pre-trained language models to create systems capable of providing expert-quality answers in various domains.
Following their lead, a diverse range of methods including DSI (Tay et al., 2022), DynamicRetriever (Zhou et al., 2023b), SEAL (Bevilacqua et al., 2022), NCI (Wang et al., 2022a), etc., have been developed, with a continuously growing body of work. These methods explore various aspects such as model training and architectures, document identifiers, incremental learning, task-specific adaptation, and generative recommendations. Figure 4 presents an overview of the GR system and we’ll provide an in-depth discussion of each associated challenge in the following sections.
3.1. Model Training and Structure
One of the core components of GR is the model training and structure, aiming to enhance the model’s ability to memorize documents in the corpus.
3.1.1. Model Training
To effectively train generative models for indexing documents, the standard approach is to train the mapping from queries to relevant DocIDs, based on standard sequence-to-sequence (seq2seq) training methods, as described in Equation (2). This method has been widely used in numerous GR research works, such as DSI (Tay et al., 2022), NCI (Wang et al., 2022a), SEAL (Bevilacqua et al., 2022), etc. Moreover, a series of works have proposed various model training methods tailored for GR tasks to further enhance GR retrieval performance, such as sampling documents or generating queries based on document content to serve as pseudo queries for data augmentation; or including training objectives for document ranking.
Specifically, DSI (Tay et al., 2022) proposed two training strategies: one is “indexing”, that is, training the model to associate document tokens with their corresponding DocIDs, where DocIDs are pre-built based on documents in corpus, which will be discussed in detail in Section 3.2; the other is “retrieval”, using labeled query-DocID pairs to fine-tune the model. Notably, DSI was the first to realize a differentiable search index based on the Transformer (Vaswani et al., 2017) structure, showing good performance in web search (Nguyen et al., 2016) and question answering (Kwiatkowski et al., 2019) scenarios. Next, a series of methods propose training methods for data augmentation and improving GR model ranking ability
Sampling Document Pieces as Pseudo Queries. In the same era, DynamicRetriever (Zhou et al., 2023b), also based on the encoder-decoder model, constructed a model-based IR system by initializing the encoder with a pre-trained BERT (Kenton and Toutanova, 2019). Besides, DynamicRetriever utilizes passages, sampled terms and N-grams to serve as pseudo queries to enhance the model’s memorization of DocIDs. Formally, the training methods can be summarized as follows:
(5) Sampled Document $:d_{s_{i}}\longrightarrow\text{DocID},i\in\{1,...,k_{d_{s}}\},$
(6) Labeled Query $:q_{i}\longrightarrow\text{DocID},i\in\{1,...,k_{q}\},$
where $d_{s_{i}}$ and $q_{i}$ denote each of the $k_{d_{s}}$ sampled document text and each of the $k_{q}$ labeled query for the corresponding DocID, respectively.
Generating Pseudo Queries from Documents. Following DSI, the NCI (Wang et al., 2022a) model was trained using a combination of labeled query-document pairs and augmented pseudo query-document pairs. Specifically, NCI proposed two strategies: one using the DocT5Query (Nogueira et al., 2019) model as a query generator, generating pseudo queries for each document in the corpus through beam search; the other strategy directly uses the document as a query, as stated in Equation (5). Similarly, DSI-QG (Zhuang et al., 2022) also proposed using a query generator to enhance training data, establishing a bridge between indexing and retrieval in DSI. This data augmentation method has been proven in subsequent works to be an effective method to enhance the model’s memorization for DocIDs, which can be expressed as follows:
(7) Pseudo $\text{Pseudo Query}:q_{s_{i}}\longrightarrow\text{DocID},i\in\{1,...,k_{q_{s}}\},$
where $q_{s_{i}}$ represents each of the $k_{q_{s}}$ generated pseudo query for the corresponding DocID.
Figure 4. A conceptual framework for a generative retrieval system, with a focus on challenges in incremental learning, identifier construction, model training and structure, and integration with downstream tasks and recommendation systems.
Improving Ranking Capability. Additionally, a series of methods focus on further optimizing the ranking capability of GR models. Chen et al. (Chen et al., 2023b) aimed at understanding the DSI model and its retrieval effectiveness, noting that the DSI model has flaws in exclusivity, completeness, and relevance ordering, proposing a multi-task distillation method to improve retrieval quality without changing the model structure, thereby obtaining better indexing and ranking capabilities. Meanwhile, LTRGR (Li et al., 2023d) proposed to directly learn how to rank paragraphs with GR models, not just generating paragraph identifiers. This framework is first trained with a sequence-to-sequence loss, then further trained with a ranking loss, allowing the model to learn ranking at the document level. Subsequently, (Zhou et al., 2023a) proposed GenRRL, which improves the overall ranking quality through reinforcement learning with relevance feedback, aligning token-level DocID generation with document-level relevance estimation after seq2seq supervised fine-tuning. Moreover,  (Li et al., 2024f) introduced DGR, which enhances GR through knowledge distillation. Specifically, DGR uses a cross-encoder as a teacher model, providing fine-grained passage ranking supervision signals, and then optimizes the GR model with a distilled RankNet loss, effectively transferring the capabilities of advanced ranking models to the GR model. ListGR (Tang et al., 2024b) also focuses on enhancing the ranking ability of GR models, by defining positional conditional probabilities, emphasizing the importance of the generation order of each DocID in the list. In addition, ListGR employs relevance calibration that adjusts the generated list of DocIDs to better align with the labeled ranking list. See Table 1 for a detailed comparison of GR methods.
3.1.2. Model Structure
Basic generative retrieval models mostly use pre-trained encoder-decoder structured generative models, such as T5 (Raffel et al., 2020) and BART (Lewis et al., 2020a), fine-tuned for the DocID generation task. To better adapt to the GR task, researchers have proposed a series of specifically designed model structures (Wang et al., 2022a; Ren et al., 2023; Lee et al., 2023b; Zhang et al., 2023e; Qiao et al., 2023; Yuan et al., 2024; Tang et al., 2024a).
Model Decoding Methods. For the semantic structured DocID proposed by DSI (Tay et al., 2022), NCI (Wang et al., 2022a) designed a Prefix-Aware Weight-Adaptive (PAWA) decoder structure. By perceiving the hierarchical prefix of DocID and adjusting the weights at different positions, this decoder can capture the semantic hierarchy of DocIDs. To overcome the capacity limitations of GR models, NP-Decoding (Lee et al., 2023b) proposed a decoding method that uses non-parametric contextualized word embeddings (as external memory) instead of traditional word embeddings as the input to the decoder. This allows the GR model not only to utilize its own parameter space but also to access and use information in contextualized word embeddings. Additionally, PAG (Zeng et al., 2024) proposed a planning-ahead generation approach, constructing DocIDs based on both sets and sequences. PAG first decodes the set-based DocID to approximate document-level scores, and then continues to decode the sequence-based DocID on this basis.
Combining Generative and Dense Retrieval Methods. Combining seq2seq generative models with dual-encoder retrieval models, MEVI (Zhang et al., 2023e) utilizes Residual Quantization (RQ) (Martinez et al., 2014) to organize documents into hierarchical clusters targeted by the generative model. This structure enables efficient retrieval of candidate clusters followed by precise document search within those clusters using Approximate Nearest Neighbor (ANN) search methods, enhancing both recall and low-latency efficiency. Similarly, Generative Dense Retrieval (GDR) (Yuan et al., 2024) merges the advantages of generative and dense retrieval in a two-step process. It first broadly matches queries to document clusters, optimizing for interaction depth and memory efficiency, and then performs precise, cluster-specific document searches, boosting both recall and scalability. This dual approach effectively combines broad matching with targeted searching, offering an optimized solution for document retrieval.
Utilizing Multiple Models. TOME (Ren et al., 2023) proposed a method that decomposes the GR task into two stages, first generating text paragraphs related to the query through an additional model, then using the GR model to generate the URL related to that paragraph, thereby improving the accuracy of model retrieval. DiffusionRet (Qiao et al., 2023) proposed a two-stage enhanced generative retrieval structure based on diffusion models. First, a sequence-to-sequence diffusion model (SeqDiffuSeq (Yuan et al., 2022)) is used to generate a pseudo-document from a query, a process similar to semantic filling of text, where the generated pseudo-document is similar to real documents in length, format, and content, rich in semantic information; secondly, another sequence-to-sequence model is used for generative retrieval based on N-gram, similar to SEAL (Bevilacqua et al., 2022) during inference, generating sets of N-grams occurring in the corpus based on the FM-Index (Ferragina and Manzini, 2000). Self-Retrieval (Tang et al., 2024a) proposed an architecture that fully integrates indexing, retrieval, and evaluation into a single large language model. By successively generating natural language indices and document segments, and performing self-evaluation to score and rank the generated documents.
Table 1. Comparisons of representative generative retrieval methods, focusing on document identifier, training data augmentation, and training objective.
Model Document Identifier Training Data Augmentation Training Objective
State Data Type Order Sample Doc Doc2Query Seq2seq DocID Ranking
GENRE (Cao et al., 2021b) Static Text Sequence - - - -
DSI (Tay et al., 2022) Static Numeric Sequence - - -
DynamicRetriever (Zhou et al., 2023b) Static Numeric Sequence - - -
SEAL (Bevilacqua et al., 2022) Static Text Sequence - - -
DSI-QG (Zhuang et al., 2022) Static Numeric Sequence - - -
NCI (Wang et al., 2022a) Static Numeric Sequence - -
Ultron (Zhou et al., 2022) Static Numeric/Text Sequence - -
CorpusBrain (Chen et al., 2022b) Static Text Sequence - - -
GenRet (Sun et al., 2023d) Learnable Numeric Sequence - -
AutoTSG (Zhang et al., 2023b) Static Text Set - - -
SE-DSI (Tang et al., 2023b) Static Text Sequence - - -
Chen et al. (Chen et al., 2023b) Static Numeric Sequence -
LLM-URL (Ziems et al., 2023) Static Text Sequence - - - - -
MINDER (Li et al., 2023e) Static Text Sequence - - -
LTRGR (Li et al., 2023d) Static Text Sequence - -
NOVO (Wang et al., 2023l) Learnable Text Set - - -
GenRRL (Zhou et al., 2023a) Static Text Sequence - -
LMIndexer (Jin et al., 2023) Learnable Numeric Sequence - -
ASI (Yang et al., 2023c) Learnable Numeric Sequence - -
RIPOR (Zeng et al., 2023) Learnable Numeric Sequence -
GLEN (Lee et al., 2023a) Learnable Text Sequence -
DGR (Li et al., 2024f) Static Text Sequence - -
ListGR (Tang et al., 2024b) Static Numeric Sequence - -
3.2. Design of Document Identifiers
Another essential component of generative retrieval is document representation, also known as document identifiers (DocIDs), which act as the target outputs for the GR model. Accurate document representations are crucial as they enable the model to more effectively memorize document information, leading to enhanced retrieval performance. Table 1 provides a detailed comparison of the states, data types, and order of DocIDs across numerous GR methods. In the following sections, we will explore the design of DocIDs from two categories: numeric-based identifiers and text-based identifiers.
3.2.1. Numeric-based Identifiers
An intuitive method to represent documents is by using a single number or a series of numbers, referred to as DocIDs. Existing methods have designed both static and learnable DocIDs.
Static DocIDs. Initially, DSI (Tay et al., 2022) introduced three numeric DocIDs to represent documents: (1) Unstructured Atomic DocID: a unique integer identifier is randomly assigned to each document, containing no structure or semantic information. (2) Naively Structured String DocID: treating random integers as divisible strings, implementing character-level DocID decoding to replace large softmax output layers. (3) Semantically Structured DocID: introducing semantic structure through hierarchical $k$-means method, allowing semantically similar documents to share prefixes in their identifiers, effectively reducing the search space. Concurrently, DynamicRetriever (Zhou et al., 2023b) also built a model-based IR system based on unstructured atomic DocID. Subsequently, Ultron (Zhou et al., 2022) encoded documents into a latent semantic space using BERT (Kenton and Toutanova, 2019), and compressed vectors into a smaller semantic space via Product Quantization (PQ) (Jégou et al., 2011; Ge et al., 2014), preserving semantic information. Each document’s PQ code serves as its semantic identifier, with additional digits added to ensure the uniqueness of each DocID if duplicates exist. MEVI clusters documents using Residual Quantization (RQ) (Martinez et al., 2014), forming semantically similar clusters for efficient search. It utilizes dual-tower and seq2seq model embeddings for a balanced performance in large-scale document retrieval.
Learnable DocIDs. Unlike previous static DocIDs, GenRet (Sun et al., 2023d) proposed learnable document representations, encoding documents into short discrete DocIDs using a discrete autoencoder. GenRet transforms documents into DocIDs through an encoder, then reconstructs documents from DocIDs using a decoder, trained to minimize reconstruction error. Furthermore, GenRet optimized the model through progressive training and diversity clustering.To ensure that DocID embeddings can reflect document content, Tied-Atomic (Nguyen and Yates, 2023) proposed to enhance DocID embeddings to reflect document content accurately. Tied-Atomic links document text with token embeddings and employs contrastive loss for DocID creation, merging generative and dense retrieval’s expressiveness and efficiency. LMIndexer (Jin et al., 2023) and ASI (Yang et al., 2023c) learned optimal DocIDs through semantic indexing, with LMIndexer using a reparameterization mechanism for unified optimization, facilitating efficient retrieval by aligning semantically similar documents under common DocIDs. ASI extends this by establishing an end-to-end retrieval framework, incorporating semantic loss functions and reparameterization to enable joint training. This approach ensures that documents with similar content are assigned similar DocIDs, enhancing retrieval accuracy and efficiency. Furthermore, RIPOR (Zeng et al., 2023) treats the GR model as a dense encoder to encode document content and using a start token for decoding. It then splits these representations into vectors via RQ (Martinez et al., 2014), creating unique DocID sequences. Furthermore, RIPOR implements a prefix-guided ranking optimization, increasing relevance scores for prefixes of pertinent DocIDs through margin decomposed pairwise loss during decoding. This ensures higher scores for relevant DocID prefixes in beam search steps, enhancing search precision.
In summary, numeric-based document representations can utilize the embeddings of dense retrievers, obtaining semantically meaningful DocID sequences through methods such as $k$-means, PQ (Jégou et al., 2011), and RQ (Martinez et al., 2014); they can also combine encoder-decoder GR models with bi-encoder DR models to achieve complementary advantages (Nguyen and Yates, 2023; Zhang et al., 2023e).
3.2.2. Text-based Identifiers
Text-based DocIDs have the inherent advantage of effectively leveraging the strong capabilities of pre-trained language models and offering better interpretability.
Document Titles. The most straightforward text-based identifier is the document title, which requires each title to uniquely represent a document in the corpus, otherwise, it would not be possible to accurately retrieve a specific document. The Wikipedia corpus used in the KILT (Petroni et al., 2021) benchmark, due to its well-regulated manual annotation, has a unique title corresponding to each document. Thus, GENRE (Cao et al., 2021b), based on the title as DocID and leveraging the generative model BART (Lewis et al., 2020a) and pre-built DocID prefix, achieved superior retrieval performance across 11 datasets in KILT. Following GENRE, GERE (Chen et al., 2022a), CorpusBrain (Chen et al., 2022b), Re3val (Song et al., 2024), and CorpusBrain++ (Guo et al., 2024) also based their work on title DocIDs for Wikipedia-based tasks. Notably, LLM-URL (Ziems et al., 2023) directly generated URLs using ChatGPT prompts, achieving commendable performance after removing invalid URLs. However, in the web search scenario (Nguyen et al., 2016), document titles in the corpus often have significant duplication and many meaningless titles, making it unfeasible to use titles alone as DocIDs. Thus, Ultron (Zhou et al., 2022) effectively addressed this issue by combining URLs and titles as DocIDs, uniquely identifying documents through keywords in web page URLs and titles.
Sub-strings of Documents. To increase the flexibility of DocIDs, SEAL (Bevilacqua et al., 2022) proposed a sub-string identifier, representing documents with any N-grams within them. Using FM-Index (a compressed full-text sub-string index) (Ferragina and Manzini, 2000), SEAL could generate N-grams present in the corpus to retrieve all documents containing those N-grams, scoring and ranking documents based on the frequency of N-grams in each document and the importance of N-grams in the query. Following SEAL, various GR models (Chen et al., 2023e; Li et al., 2023e, d, 2024f) also utilized sub-string DocIDs and FM-Index during inference. For a more comprehensive representation of documents, MINDER (Li et al., 2023e) proposed multi-view identifiers, including generated pseudo queries from document content via DocT5Query (Nogueira et al., 2019), titles, and sub-strings. However, compared to single DocID methods, multi-DocID GR methods required more memory use and inference time. This multi-DocID approach was also used in LTRGR (Li et al., 2023d) and DGR (Li et al., 2024f). Concurrently, SE-DSI (Tang et al., 2023b) also demonstrated the effectiveness of using pseudo queries alone as DocIDs.
Term Sets. Unlike the sequential DocIDs described earlier, AutoTSG (Zhang et al., 2023b) proposed a term set-based document representation, using keywords extracted from titles and content, rather than predefined sequences, allowing for retrieval of the target document as long as the generated term set is included in the document’s identifiers. During inference, AutoTSG designed a constrained greedy search to ensure each term generated stepwise was an effective identifier belonging to some document. Recently, PAG (Zeng et al., 2024) also constructed DocIDs based on sets of key terms, disregarding the order of terms, which allows for document-level scoring to be approximated in decoding.
Learnable DocIDs. Text-based identifiers can also be learnable. Similarly based on term-sets, NOVO (Wang et al., 2023l) proposed learnable continuous N-grams constituting term-set DocIDs. Through the Denoising Query Modeling (DQM) task, the model learned to generate queries from documents with noise, thereby implicitly learning to filter out document N-grams more relevant to queries. NOVO also utilized the retrieval task to update the semantic representations of document identifiers, i.e., improving the document’s semantic representation by updating N-gram embeddings. Later, GLEN (Lee et al., 2023a) designs dynamic lexical DocIDs and is trained through a two-phase index learning strategy. Firstly, the keyword-based DocID assignment phase, GLEN defines DocIDs by extracting keywords from documents using self-supervised signals and learns them. Secondly, the ranking-based DocID refinement phase, GLEN learns dynamic DocIDs by directly integrating query-document relevance through two loss functions. During inference, GLEN utilizes collision-free inference, ranking documents using DocID weights without incurring additional overhead.
3.3. Incremental Learning on Dynamic Corpora
Prior studies have focused on generative retrieval from static document corpora. However, in reality, the documents available for retrieval are continuously updated and expanded. To address this challenge, researchers have developed a range of methods to optimize GR models for adapting to dynamic corpora.
Optimizer and Document Rehearsal. At first, DSI++ (Mehta et al., 2022) aims to address the incremental learning challenges within dynamic document corpora encountered by DSI (Tay et al., 2022). DSI++ modifies the training by optimizing flat loss basins through the Sharpness-Aware Minimization (SAM) optimizer, thereby stabilizing the learning process of the model. It also introduces generative memory by employing DocT5Query (Nogueira et al., 2019) to generate pseudo queries for documents in the existing corpus as training data augmentation, mitigating the forgetting issue of GR models.
Constrained Optimization Addressing the scenario of real-time addition of new documents, such as news or scientific literature IR systems, IncDSI (Kishore et al., 2023) views the addition of new documents as a constrained optimization problem to find optimal representations for the new documents. This approach aims to (1) ensure new documents can be correctly retrieved by their relevant queries, and (2) maintain the retrieval performance of existing documents unaffected. IncDSI manages to add each document within approximately 20-50ms, significantly reducing the time and computational resources required compared to full model retraining, while maintaining competitive retrieval performance.
Incremental Product Quantization. CLEVER (Chen et al., 2023d), based on Product Quantization (PQ) (Jégou et al., 2011), proposes Incremental Product Quantization (IPQ) for generating PQ codes as DocIDs for documents. Compared to traditional PQ methods, IPQ designs two adaptive thresholds to update only a subset of centroids instead of all, maintaining the indices of updated centroids constant. This method reduces computational costs and allows the system to adapt flexibly to new documents. To mitigate forgetting, CLEVER employs a memory-enhanced learning mechanism, maintaining a dynamic memory bank to store example documents similar to the new ones.
Fine-tuning Adatpers for Specific Tasks. CorpusBrain++ (Guo et al., 2024) introduces the KILT++ benchmark for continuously updated KILT (Petroni et al., 2021) tasks and designs a dynamic architecture paradigm with a backbone-adapter structure. By fixing a shared backbone model to provide basic retrieval capabilities while introducing task-specific adapters to incrementally learn new documents for each task, it effectively avoids catastrophic forgetting. During training, CorpusBrain++ generates pseudo queries for new document sets and continues to pre-train adapters for specific tasks. Moreover, it employs document clustering based on semantic similarity and a retraining strategy to maintain memory of older documents by revisiting them.
3.4. Downstream Task Adaption
Generative retrieval methods, apart from addressing retrieval tasks individually, have been tailored to various downstream generative tasks. These include fact verification (Thorne et al., 2018), entity linking (Hoffart et al., 2011), open-domain QA (Kwiatkowski et al., 2019), dialogue (Dinan et al., 2018), slot filling (Levy et al., 2017), among others, as well as knowledge-intensive tasks (Petroni et al., 2021), code (Lu et al., 2021), conversational QA (Adlakha et al., 2022), and multi-modal retrieval scenarios (Lin et al., 2014), demonstrating superior performance and efficiency. These methods are discussed in terms of separate training, joint training, and multi-modal generative retrieval.
3.4.1. Separate Training
For fact verification tasks (Thorne et al., 2018), which involve determining the correctness of input claims, GERE (Chen et al., 2022a) proposed using an encoder-decoder-based GR model to replace traditional indexing-based methods. Specifically, GERE first utilizes a claim encoder to encode input claims, and then generates document titles related to the claim through a title decoder to obtain candidate sentences for corresponding documents. Finally, an evidence decoder generates evidence sentence identifiers, resulting in improvements in time, memory consumption, and performance.
Knowledge-Intensive Language Tasks. For Knowledge-Intensive Language Tasks (KILT) (Petroni et al., 2021), CorpusBrain (Chen et al., 2022b) introduced three pre-training tasks to enhance the model’s understanding of query-document relationships at various granularities: Internal Sentence Selection, Leading Paragraph Selection, and Hyperlink Identifier Prediction. Similarly, UGR (Chen et al., 2023e) proposed using different granularities of N-gram DocIDs to adapt to various downstream tasks, unifying different retrieval tasks into a single generative form. UGR achieves this by letting the GR model learn prompts specific to tasks, generating corresponding document, passage, sentence, or entity identifiers.
Futhermore, DearDR (Thorne, 2022) is an efficient data-centric self-regressive document retrieval method. It utilizes remote supervision and self-supervised learning techniques, using Wikipedia page titles and hyperlinks as training data. The model samples sentences from Wikipedia documents as input and trains a self-regressive model to decode page titles or hyperlinks, or both, without the need for manually labeled data. The model achieves zero-shot implementation of Wikipedia-based fact verification tasks and further optimizes performance through fine-tuning. Re3val (Song et al., 2024) proposes a retrieval framework combining generative reordering and reinforcement learning. It first reorders retrieved page titles using context information obtained from a dense retriever, then optimizes the reordering using the REINFORCE algorithm to maximize rewards generated by constrained decoding. By improving page title reordering and context selection, Re3val achieves more accurate information retrieval.
Multi-hop retrieval. In multi-hop retrieval tasks, which require iterative document retrievals to gather adequate evidence for answering a query, GMR (Lee et al., 2022c) proposed to employ language model memory and multi-hop memory to train a generative retrieval model, enabling it to memorize the target corpus and simulate real retrieval scenarios through constructing pseudo multi-hop query data, achieving dynamic stopping and efficient performance in multi-hop retrieval tasks.
Code Retrieval. CodeDSI (Nadeem et al., 2022) is an end-to-end generative code search method that directly maps queries to pre-stored code samples’ DocIDs instead of generating new code. Similar to DSI (Tay et al., 2022), it includes indexing and retrieval stages, learning to map code samples and real queries to their respective DocIDs. CodeDSI explores different DocID representation strategies, including direct and clustered representation, as well as numerical and character representations, showing superior performance compared to traditional methods on 1K and 10K scales, with numerical DocIDs performing better than alphabetic ones.
Conversational Question Answering. GCoQA (Li et al., 2023c) is a generative retrieval method for conversational QA systems. Unlike traditional methods relying on bi-encoder architecture and similarity matching, GCoQA directly generates DocIDs for passage retrieval. This method focuses on key information in the dialogue context at each decoding step, achieving more precise and efficient passage retrieval and answer generation, thereby improving retrieval performance and overall system efficiency.
3.4.2. Joint Training
The methods in the previous section involve separately training generative retrievers and downstream task generators. However, due to the inherent nature of GR models as generative models, a natural advantage lies in their ability to be jointly trained with downstream generators to obtain a unified model for retrieval and generation tasks.
Multi-decoder Structure. UniGen (Li et al., 2024g) proposes a unified generation framework to integrate retrieval and question answering tasks, bridging the gap between query input and generation targets using connectors generated by large language models. UniGen employs shared encoders and task-specific decoders for retrieval and question answering, introducing iterative enhancement strategies to continuously improve the performance of both tasks. It demonstrates superior performance on both web search and question answering tasks.
Multi-task Training. Later, CorpusLM (Li et al., 2024a) introduces a unified language model that integrates GR, closed-book generation, and retrieval-augmented generation to handle various knowledge-intensive tasks. The model adopts a multi-task learning approach and introduces ranking-guided DocID decoding strategies and continuous generation strategies to improve retrieval and generation performance. In addition, CorpusLM designs a series of auxiliary DocID understanding tasks to deepen the model’s understanding of DocID semantics. Experimental results validate the effectiveness and potential of CorpusLM (T5 (Raffel et al., 2020) and Llama2 (Touvron et al., 2023b) variants) for knowledge-intensive language tasks.
3.4.3. Multi-modal Generative Retrieval
Generative retrieval methods can also leverage multi-modal data such as text, images, etc., to achieve end-to-end multi-modal retrieval.
Tokenizing Images to DocID Sequences. At first, IRGen (Zhang et al., 2023g) transforms image retrieval problems into generative problems, predicting relevant discrete visual tokens, i.e., image identifiers, through a seq2seq model given a query image. Its core innovation lies in its semantic image tokenizer, which converts global image features into short sequences capturing high-level semantic information. Unlike traditional methods that handle feature extraction and ANN search separately, IRGen achieves end-to-end differentiable search, optimizing directly from the final retrieval target, thereby enhancing retrieval accuracy and efficiency.
Advanced Model Training and Structure. Later, GeMKR (Long et al., 2024) combines LLMs’ generation capabilities with visual-text features, designing a generative knowledge retrieval framework. It first guides multi-granularity visual learning using object-aware prefix tuning techniques to align visual features with LLMs’ text feature space, achieving cross-modal interaction. GeMKR then employs a two-step retrieval process: generating knowledge clues closely related to the query and then retrieving corresponding documents based on these clues. It aims to improve knowledge retrieval efficiency and accuracy in multi-modal scenarios. GRACE (Long et al., 2024) achieves generative cross-modal retrieval method by assigning unique identifier strings to images and training multi-modal large language models (MLLMs) (Awadalla et al., 2023) to memorize the association between images and their identifiers. The training process includes (1) learning to memorize images and their corresponding identifiers, and (2) learning to generate the target image identifiers from textual queries. GRACE explores various types of image identifiers, including strings, numbers, semantics, identifiers, and atomic identifiers, to adapt to different memory and retrieval requirements.
3.5. Generative Recommender Systems
Recommendation systems, as an integral part of the information retrieval domain, are currently undergoing a paradigm shift from discriminative models to generative models. Generative recommendation systems do not require the computation of ranking scores for each item followed by database indexing, but instead accomplish item recommendations through the direct generation of IDs. In this section, several seminal works, including P5 (Geng et al., 2022), GPT4Rec (Li et al., 2023h), TIGER (Rajput et al., 2023), SEATER (Si et al., 2023), IDGenRec (Tan et al., 2024b), LC-Rec (Zheng et al., 2023) and ColaRec (Wang et al., 2024b), are summarized to outline the development trends in generative recommendations.
Numeric-based Item IDs. P5 (Geng et al., 2022) transforms various recommendation tasks into different natural language sequences, designing a universal, shared framework for recommendation completion. This method, by setting unique training objectives, prompts, and prediction paradigms for each recommendation domain’s downstream tasks, serves well as a backbone model, accomplishing various recommendation tasks through generated text. This approach demonstrates the viability and flexibility of generative models in recommendation systems. In generative retrieval, effective indexing identifiers have been proven to significantly enhance the performance of generative methods. Similarly, TIGER (Rajput et al., 2023) initially learns a residual quantized autoencoder to generate semantically informative indexing identifiers for different items. It then trains a transformer-based encoder-decoder model with this semantically informative indexing identifier sequence to generate item identifiers for recommending the next item based on historical sequences.
Focusing solely on semantic information and overlooking the collaborative filtering information under the recommendation context might limit the further development of generative models. Therefore, after generating semantic indexing identifiers similar to TIGER using a residual quantized autoencoder with uniform semantic mapping, LC-Rec (Zheng et al., 2023) also engages in a series of alignment tasks, including sequential item prediction, explicit index-language alignment, and recommendation-oriented implicit alignment. Based on the learned item identifiers, it integrates semantic and collaborative information, enabling large language models to better adapt to sequence recommendation tasks.
Text-based Item IDs. IDGenRec (Tan et al., 2024b) innovatively combines generative recommendation systems with large language models by using human language tokens to generate unique, concise, semantically rich and platform-agnostic texual identifiers for recommended items. The framework includes a text ID generator trained on item metadata with a diversified ID generation algorithm, and an alternating training strategy that optimizes both the ID generator and the LLM-based recommendation model for improved performance and accuracy in sequential recommendations. The zero-shot performance of IDGenRec is comparable to, or even surpasses, certain traditional recommendation models that rely on supervised training, highlighting its potential as a foundational model for recommendation systems. SEATER (Si et al., 2023) designs a balanced k-ary tree-structured indexes, using a constrained k-means clustering method to recursively cluster vectors encoded from item texts, obtaining equal-length identifiers. Compared to the method proposed by DSI (Tay et al., 2022), this balanced k-ary tree index maintains semantic consistency at every level. It then trains a Transformer-based encoder-decoder model and enhances the semantics of each level of indexing through contrastive learning and multi-task learning. ColaRec (Wang et al., 2024b) integrates collaborative filtering signals and content information by deriving generative item identifiers from a pretrained recommendation model and representing users via aggregated item content. Then it uses an item indexing generation loss and contrastive loss to align content-based semantic spaces with collaborative interaction spaces, enhancing the model’s ability to recommend items in an end-to-end framework.
4. Reliable Response Generation: Direct Information Accessing with Generative Language Models
Figure 5. An Illustration of strategies for enhancing language models to generate user-centric and reliable responses, including model internal knowledge memorization and external knowledge augmentation.
The rapid advancement of large language models has positioned them as a novel form of IR system, capable of generating reliable responses directly aligned with users’ informational needs. This not only saves the time users would otherwise spend on collecting and integrating information but also provides personalized, user-centric answers tailored to individual users.
However, challenges remain in creating a grounded system that delivers faithful answers, such as hallucination, prolonged inference time, and high operational costs. This section will discuss strategies for constructing a faithful GenIR system, focusing on both optimizing the model internally and enhancing it with external knowledge.
4.1. Internal Knowledge Memorization
To develop an user-friendly and reliable IR system, the generative model should be equipped with comprehensive internal knowledge. Optimization of the backbone generative model can be categorized into three aspects: structural enhancements, training strategies, and inference techniques.
4.1.1. Model Structure
With the advent of generative models, a variety of methods have been introduced to improve model structure and enhance generative reliability. We aim to discuss the crucial technologies contributing to this advancement in this subsection.
(1) Model Scaling Model parameter scaling is a pivotal factor influencing performance. Contemporary language models predominantly employ the Transformer architecture, which has been observed that scaling both the model parameters and the training data enhances the model’s capacity to retain knowledge and capabilities (Kaplan et al., 2020). For instance, in the GPT (Radford et al., 2018, 2019; Brown et al., 2020; Achiam et al., 2023) series and LLaMA (Touvron et al., 2023a, b) family, models with larger parameter sizes tend to perform better on diverse downstream tasks. These include few-shot learning, language understanding, and generation (Chowdhery et al., 2023). In addition, scaling the model contributes to improved instruction-following capabilities (Qin et al., 2024), enabling a more adept comprehension of user intent and generating responses that better satisfy user requests.
(2) Model integration Model integration is an effective and intuitive method to enhance the reliability of generated outputs by capitalizing on diverse strengths inherent in various models. This diversity arises from differences in training data and architectural frameworks. The predominant approach to model integration is the ’Mixture of Experts’ (MoE) (Jacobs et al., 1991), which utilizes a gating mechanism to selectively activate sections of network parameters during inference. MoE facilitates the amalgamation of the strengths of multiple expert models, greatly increasing the effective parameters of the model without inflating inference costs. Consequently, this leads to enhanced performance (Fedus et al., 2022; Lepikhin et al., 2020; Du et al., 2021; Jiang et al., 2024). This method also boasts impressive scalability: the overall efficacy of the MoE model is augmented in tandem with the expanding parameter volume and the number of expert models (Clark et al., 2022).
In contrast to the MoE mechanism that involves training multiple expert models and a gating system from scratch, the LLM-Blender framework (Jiang et al., 2023a) employs a ranker and a fuser module to organize and merge answers from diverse LLMs. This approach is applicable to a wide array of large-scale models, including black-box types; however, it presents challenges associated with high deployment costs.
4.1.2. Training and Inference
In the model training stage, methods to enhance the reliability of answers can be categorized into two aspects: training data optimization and training methods optimization.
(1) Training Data Optimization The quality of training data substantially affects the reliability of model outputs. Noise, misinformation, and incomplete information can disrupt the learning process of models, leading to hallucation and other issues. Many studies have focused on the processing of model training data. (Gunasekar et al., 2023) used GPT3.5 to artificially create textbooks filled with examples and language descriptions as training data. They minorly fine-tuned the model with a small amount of data, resulting in significant improvements on downstream tasks. LIMA (Zhou et al., 2024b) used dialogues from community forums to construct a small-scale fine-tuning dataset, enhancing the model’s conversation capabilities during the alignment phase.
Concerning the fact that a large part of training data comes from automatically crawled internet page data, which contains many redundancies, some works have attempted to deduplicate the training dataset. Using a combination of suffix array (Manber and Myers, 1990) and MinHash (Broder, 1997) techniques, Lee et al. (Lee et al., 2022a) designed an approximate matching algorithm to identify redundant content originating from the same source, hence reducing the proportion of answers the model directly reproduces from training data.
(2) Training Methods Optimization In addition to conventional training methods, additional techniques have been proposed to improve the factuality of model outputs. Given that the autoregressive training objective of language models can lead to blind imitation of training data, MixCL (Sun et al., 2022a) incorporated contrastive learning into the training objective of language models. It used an external knowledge base to identify correct knowledge snippets, incorporated contrastive learning loss function to reduce the generation probability of tokens in incorrect knowledge snippets, thereby enhancing model reliability. CaliNet (Dong et al., 2022) utilized a contrastive method to assess erroneous knowledge learned by the model and fine-tuned the parameters of the FFN layer to rectify these errors. FactTune (Ovadia et al., 2023a) incorporated factuality assessment during the RLHF phase. It used several automatic evaluation methods, including Factscore (Min et al., 2023), to rank the factuality of model outputs. Consequently, it utilized the DPO (Rafailov et al., 2023) method to teach the model factuality preference ranking.
Apart from enhancing the internal knowledge reliability of the model during the training phase, the inference stage significantly impacts the reliability of answers. The overall inference process consists of two parts: the inputting of corresponding instructions by the user to the model and the decoding of the corresponding response tokens by the model. The approach to increase generation reliability during the inference stage can also be divided into two parts: prompting engineering and decoding strategy.
(3) Prompt Engineering Prompting method plays a vital role in guiding the model. A well-designed prompt can better promote the model’s internal capabilities to provide more accurate answers. The Chain-of-Thought (CoT) (Wei et al., 2022b) prompting method guides the model to explicitly decompose the question into a reasoning chain during the decoding process and to reason out the answer step by step according to the reasoning chain. This method can improve the response accuracy for each sub-question, grounding the final question on the most accurate intermediate step, thus enhancing the reliability of the answer. Further, CoT-SC (Wang et al., 2023g) builds upon this by sampling multiple answers and choosing the most consistent one as the final answer. The Tree of Thoughts (Yao et al., 2023b) expands the single reasoning path of CoT to multiple paths and synthesizes the reasoning outcomes of these paths to arrive at the final answer. The Chain-of-Verification (CoVE) (Dhuliawala et al., 2023) introduces a self-reflection mechanism within the prompts of LLM. The prompt causes the LLM to generate a draft response, and then formulates a validation plan for each statement within the response to check for factual inaccuracies. If errors are found, corrections are carried out, effectively enhancing the factual accuracy of the response.
Besides, some prompting methods enhance the grounding of answers on more accurate content by prompting the model to output relevant internal knowledge. For example, RECITE (Sun et al., 2022b) and GenRead (Yu et al., 2022) utilize a sampling method to prompt the model to output certain relevant knowledge fragments, which are then used to bolster the question-answering process.
(4) Decoding Strategy Decoding strategies are another critical factor influencing the reliability of model-generated responses. An appropriate decoding method can maintain the reliability and diversity of a model’s response. In common gready search or beam search approaches, models may produce responses that are high in general applicability but lack diversity, compromising the quality of the generated content. Nucleus sampling (Lee et al., 2022b) is proposed to tackle this issue. By sampling within a set probability range for tokens, it can ensure better diversity for each generated token, and achieve results that balance variety and reliability. Building on this, Lee et al. introduced Factual-Nucleus Sampling (Lee et al., 2023c), which employs a dynamic threshold that decays for subsequent tokens, ensuring that later parts of a generated sentence are not adversely affected by earlier parts that may lack factual content. Wan et al. (Wan et al., 2023) proposed a faithfulness-aware decoding method to enhance the faithfulness of the beam-search approach. This method incorporates a Ranker to reorder the generated beam sequences and a lookahead method to avoid selecting tokens that may lead to unfaithfulness in further generation, thereby improving the faithfulness of the results.
Apart from directly modifying the decoding method, numerous studies have influenced the model’s decoding distribution by utilizing or modifying the information in the hidden layers. DoLa (Chuang et al., 2023) uses the distributional differences between the hidden layer and the output layer as the distribution for the next token, capturing factual knowledge or key terms that are newly learned by the output layer relative to the base network, hence increasing the likelihood of generating these terms. Inference-Time Intervention (ITI) (Li et al., 2023b) focuses on identifying attention heads highly correlated with response correctness and adjusting their orientations, then moderating the activation of these attention heads. This method achieves more truthful generation with minimal interference to the model.
In practical applications, prompts often contain additional textual content to provide extra information, which may conflict with the model’s inherent knowledge, leading to hallucinations. Shi et al. (Shi et al., 2023a) proposed the CAD method, which compares the output distributions before and after the addition of extra information. This reduces the weight given to the model’s own knowledge, thus avoiding conflicts between the two that might result in factual inaccuracies.
4.1.3. Knowledge Updating
In real-life scenarios, information is constantly evolving, and therefore, the GenIR system needs to continuously acquire the latest knowledge to meet users’ information needs. Since the model’s knowledge storage is limited, knowledge updating is necessary to ensure more reliable generated responses. In this section, we will discuss existing methods for knowledge updating from two perspectives: incremental learning and knowledge editing.
(1) Incremental Learning. Incremental learning refers to the ability of machine learning models to continuously learn new skills and tasks while retaining previously acquired knowledge (Wang et al., 2023k; Wu et al., 2024a; Wang et al., 2023m; Zhang et al., 2024a). In the GenIR system, it is crucial to enable the language model to memorize the latest information while preventing the forgetting of previous knowledge.
One approach is Incremental Pre-training, which does not rely on supervised data but continues pre-training on continuously updated corpora to alleviate catastrophic forgetting. For example, Baidu proposed the ERNIE 2.0 framework (Sun et al., 2020), which enhances language understanding through continuous multi-task learning. Jang et al. (Jang et al., 2021) introduced the concept of Continual Knowledge Learning (CKL) to explore how LLMs can update and retain knowledge in the face of rapidly changing world knowledge. They also created new benchmarks and evaluation metrics such as FUAR. Cossu et al. (Cossu et al., 2022) studied continual pre-training for language and vision and found that self-supervised or unsupervised pre-training methods are more effective in retaining previous knowledge compared to supervised learning. Additionally, Ke et al. (Ke et al., 2023) proposed Domain Adaptive Pre-training (DAP-training) to improve the model’s adaptability to new domains while preventing the forgetting of previously learned knowledge using techniques like soft masking and contrastive learning. For domain-specific model construction, Xie et al. (Xie et al., 2023) introduced FinPythia-6.9B, an efficient continual pre-training method specifically designed for constructing large-scale language models in the financial domain.
On the other hand, Incremental Fine-tuning involves training the model using only labeled data. For instance, Progressive Prompts (Razdaibiedina et al., 2023) creates new soft prompts for each new task and concatenates them in the order of their occurrence, effectively transferring knowledge and mitigating forgetting. DynaInst (Mok et al., 2023) improves the lifelong learning performance of pre-trained language models using parameter regularization and experience replay. It employs dynamic instance selection and task selection mechanisms to optimize learning efficiency under resource-constrained settings. Jang et al. (Jang et al., 2023b) challenge traditional multi-task prompt fine-tuning methods by fine-tuning expert language models on individual tasks. Suhr et al. (Suhr and Artzi, 2024) adopt a feedback-driven continual learning approach applicable to instruction-following agents. In human-agent interaction, users control the agent using natural language and convert feedback into immediate rewards through contextual bandits, further optimizing the learning process. Furthermore, O-LoRA (Wang et al., 2023a) demonstrates its superiority in the field of continual learning by learning new tasks in low-rank subspaces while maintaining orthogonality between these subspaces, significantly reducing interference between tasks. Peng et al. (Peng et al., 2024) propose a scalable language model that dynamically adjusts parameters based on different task requirements, effectively reducing the forgetting of previous task knowledge during new task learning.
(2) Knowledge Editing. Knowledge editing refers to the process of modifying and updating existing knowledge within language models (Wang et al., 2023m; Mazzia et al., 2023), which is distinct from incremental learning that focuses on adapting to new domains or tasks. Knowledge editing methods, by editing the weights or layers of a model, can correct erroneous facts and incorporate new knowledge. This makes it an important technology before deploying GenIR systems. There are primarily three paradigms for internal knowledge editing within language models: adding trainable parameters, locate-then-edit, and meta-learning.
One method of Adding Trainable Parameters is by integrating new single neurons (patches) in the final feed-forward neural network (FFN) layer, such as T-Patcher (Huang et al., 2023c) and CaliNET (Dong et al., 2022). These neurons serve as trainable parameters that introduce new errors and adjust the model’s behavior. Alternatively, discrete code-book modules are introduced in the middle layers of the language model, as in GRACE (Hartvigsen et al., 2023), to adjust and correct information.
Moerover, the Locate-then-Edit method first identifies the parameters corresponding to specific knowledge and then updates these targeted parameters directly. Common techniques involve identifying key-value pairs in the FFN matrix, known as ”knowledge neurons,” and updating them (Dai et al., 2022). Techniques like ROME (Meng et al., 2022) use causal mediation analysis to pinpoint areas needing editing, and MEMIT (Meng et al., 2023) builds on ROME to implement synchronized editing in various scenarios. Moreover, methods such as PMET (Li et al., 2024c) employ attention mechanisms for editing, while BIRD (Ma et al., 2023b) has introduced a bidirectional inverse relation modeling approach.
Meta Learning, which is another paradigm, uses hyper-networks to generate the necessary updates for model editing. KE (Knowledge Editor) (Cao et al., 2021a) predicts weight updates for each data point using a hyper-network. MEND (Mitchell et al., 2022), by taking low-order decomposition of gradients as input, learns to rapidly edit language models to enhance performance. Additionally, MALMEN (Tan et al., 2023) manages to separate the computations of hyper-networks and language models, facilitating the editing of multiple facts under a limited memory budget. These meta-learning mechanisms enable models to swiftly adapt to new knowledge and tasks and perform the necessary edits.
4.2. External Knowledge Augmentation
Although large language models have demonstrated significant effectiveness in response generation, issues such as susceptibility to hallucinations, difficulty to handle in-domain knowledge and challenges with knowledge updating persist. Augmenting the model’s generative process with external knowledge sources can serves as an effectively address to tackle these issues. Based on the form of external knowledge employed, these approaches can be classified into retrieval augmentation and tool augmentation.
4.2.1. Retrieval Augmentation
Retrieval-Augmented Generation (RAG) enhance the response of generative models by combining them with a retrieval mechanism (Lewis et al., 2020b; Izacard and Grave, 2020; Zhou et al., 2024a). By querying a large collection of documents, information that is relevant to the input query can be fetched and integrated into input of the generative model. RAG enables generative models to be grounded in existing reliable knowledge, significantly improving the reliability of model generation. Typically, a RAG method involves a retriever and a generator. Based on the interaction flow between these two, RAG methods can be divided into four categories (Gao et al., 2024).
(1) Sequential RAG: Sequential RAG operates on a linear progression, where the retriever first retrieves relevant information and the generator utilizes information to directly complete the response generation process.
The basic form of sequential RAG is a ”Retrieve-Read” framework (Ma et al., 2023a). Early works complete the response generation process through joint training of retriever and generator (Lewis et al., 2020b; Guu et al., 2020; Borgeaud et al., 2022) or separate training (Izacard and Grave, 2020). Although it has achieved good results, there is a need for pre-training of the language model, which is expensive and not conducive to generalization. The In-Context RALM (Ram et al., 2023) solves this problem by directly using the retrieval document as input, utilizing the model’s in-context learning capability to understand the document without any training.
With the widespread adoption of large language models, most subsequent works is built on the foundation of a frozen generator. AAR method (Yu et al., 2023) fine-tuned a general retriever to adapt to the information acquisition preferences of the generative model. LLM-embedder (Zhang et al., 2023f) using reward produced by LLM to train an embedding model dedicated to retrieval augmentation. ARL2 (Zhang et al., 2024b) leverages LLM to annotate relevance score in training set, and train a retriever using contrastive learning. Several works introduce pre-retrieval and post-retrieval process (Gao et al., 2024) into the sequential pipeline to enhance the overall efficiency. In pre-retrieval process, the RRR model (Ma et al., 2023a) introduces a rewriter module before the retriever. The rewriter is trained by utilizing generator’s feedback and can enable the retrieval system to provide more suitable information for generation.
In post-retrieval process, information compressors are proposed to filter out irrelevant content from documents, avoiding misleading generator’s response. (Liu et al., 2023b; Cuconasu et al., 2024; Jin et al., 2024) RECOMP (Xu et al., 2023b) introduce both abstractive and extractive compressors to generate concise summary for retrieved documents. LLMLingua (Jiang et al., 2023d) calculates the importance of each token based on the perplexity provided by the generative model, while retaining important tokens. Furthermore, LongLLMLingua (Jiang et al., 2023c)introduces query-aware compression on its basis and rerank the retrieved documents based on the calculated importance score to alleviate the loss in the middle phenomenon (Liu et al., 2023b). PRCA (Yang et al., 2023a) employs reinforcement learning to train a text compressor adaptable to black-box LLMs and various retrievers, functioning as a versatile plug-in for multiple scenarios.
(2) Branching RAG: In Branching RAG framework, the input query is processed across multiple pipelines, and each pipeline may involve the entire process in the sequential pipeline. The outputs from all pipelines merged to form the final response. Compared to the sequential RAG process, this approach allows for finer-grained handling of the query or retrieval results.
In pre-retrieval stage, TOC (Kim et al., 2023) applies few-shot prompting to recursively decompose complex, ambiguous questions into clear, disambiguated sub-questions. These sub-questions are structured in a tree structure, with relevant documents retrieved for each. Using all valid question nodes, a long-form answer covering all sub-questions is generated. BlendFilter (Wang et al., 2024d) leverages prompts to enhance the original query through both internal and external knowledge. The augmented queries are then used to retrieve related documents, which are subsequently merged. When facing complex questions from users, this type of method can break down the questions and then answer them, improving the comprehensiveness of the response.
In post-retrieval stage, REPLUG (Shi et al., 2023b) inputs each retrieved document along with the query into a generator to obtain a predicted probability distribution. These individual distributions are then amalgamated to yield the ultimate predictive probability distribution. GenRead (Yu et al., 2022) prompts LLM to generate related documents on its own while utilizing the retriever to retrieve documents. These two sets of documents are then merged as input. The versatility in designing prompts for LLM to create related documents allows for a broader coverage of content, thus enhancing the likelihood of meeting the user’s query requirements more effectively.
(3) Conditional RAG: The Conditional RAG framework adapts to various query types through distinct processes, improving the system’s flexibility. Since there can be knowledge conflict between the knowledge from retrieved documents and the generator’s own knowledge, RAG’s effectiveness isn’t consistent across all scenarios. To address this, common conditional RAG methods include a decision-making module that determines whether to engage the retrieval process for each specific query. SKR (Wang et al., 2023d) builds a binary classification model trained on a dataset compiled from questions that LLMs can and cannot answer. This model is leveraged in inference stage to discern whether a given query should utilize retrieval. Labeling for the training data is gathered by directly prompting the model to ascertain if external knowledge is required for generation. Self-DC (Wang et al., 2024c) employs the confidence score of the model’s response to determine the necessity of retrieval. Based on the confidence score, queries are categorized into three groups: unknown, uncertainty, and known. Queries deemed unknown are processed through a sequential RAG pipeline, while those with uncertainty are broken down into sub-questions to generate answers. Rowen (Ding et al., 2024) introduces a multilingual detection module that transforms the original question into semantically equivalent perturbed questions and collects their responses. The decision to retrieve is then based on measuring the consistency across these responses.
(4) Loop RAG: Loop RAG involves deep interactions between the retriever and generator components. Owing to multi-turn retrieval and generation processes, accompanied by comprehensive interactions, it excels at handling complex and diverse input queries, yielding superior results in response generation.
ITER-RETGEN (Shao et al., 2023) introduce an iterative retrieval and generation framework. For a given query, it initially performs retrieval-augmented generation, retrieving relevant documents and then generate an answer. Subsequently, it undertakes generation-augmented retrieval, where it continues to retrieve based on the content generated in the previous step, and synthesizes text using the newly retrieved documents. This process of alternating the two steps is repeated a fixed number of times to produce the final answer. IR-COT (Trivedi et al., 2022) has a similar overall procedure to ITER-RETGEN, but its iteration pause is contingent on the model’s own generative process.
FLARE (Jiang et al., 2023e) introduces a strategy of concurrent retrieval as responses are being generated, in contrast to conducting retrieval only after a complete response has been produced. For each new sentence generated, the framework evaluates the need for retrieval based on the LLM’s confidence score for that sentence. If needed, it formulates a query based on the sentence to retrieve relevant information, and then regenerates the sentence. This method dynamically supplements the LLM with the necessary information, enhancing the reliability of the generated content. COG (Lan et al., 2023) models the generation process as a continual retrieval of segments from an external corpus and subsequent copying. The generator’s primary role therein is to produce certain conjunction words to preserve sentence fluency. Self-RAG (Asai et al., 2023) adds special tokens into the vocabulary to allow the generator to decide whether to retrieve, the importance of the retrieved document, and whether to perform a critique.
Some works focus on deconstructing complex inquiries into sub-questions, addressing these individually to produce a more dependable response. (Press et al., 2023) guides LLM to decompose complex questions into sub-questions, responds to each individually using retrieved results, and ultimately synthesizes the answers to all sub-questions to form the final response. Building upon this, RET-Robust (Yoran et al., 2023) incorporates a NLI model to evaluate whether the retrieved documents can substantiate the answers to sub-questions, thereby minimizing the risk of the LLM being misled by irrelevant information.
4.2.2. Tool Augmentation
Although retrieval-augmented techniques have significantly improved upon the blind spots of generator’s self-knowledge, these methods struggle with the rapid and flexible update of information since they rely on the existence of information within an external corpus of documents. Tool augmentation, on the other hand, excels in addressing this issue by invoking various tools that allow for the timely acquisition and usage of the latest data, including finance, news, and more. Moreover, tool augmentation expands the scope of responses a model can offer, such as language translation, image generation, and other tasks, to more comprehensively meet users’ information retrieval needs. There are four categories of tools that can be utilized to construct a more reliable information retrieval system:
(1) Search Engine: Common search engine tools like Google Search and Bing Search help answer frequent and time-sensitive queries effectively. Self-Ask (Press et al., 2023) initially decomposes complex questions into multiple sub-questions, then uses search engine to answer each sub-question, and finally generating a comprehensive answer to the complex question. ReAct (Yao et al., 2022) embeds search engine calls into the model’s reasoning process, allowing the generative model to determine when to make calls and what queries to input for more flexible reasoning. New Bing can automatically search relevant information from Bing based on user input, yielding reliable and detailed answers, including citation annotations in the generated content.
Some works have also built advanced conversational systems based on tools like search engines. Internet-Augmented Generation (Komeili et al., 2022) enhances the quality of conversational replies by using search engine during conversations. LaMDA (Thoppilan et al., 2022) and BlenderBot (Shuster et al., 2022) combine search engines with conversational agents, constantly accessing internet information to enrich conversation factualness. WebGPT (Nakano et al., 2021) and WebCPM (Qin et al., 2023a) directly teache models to perform human-like browser operations by generating commands such as Search, Click, and Quote, facilitating the automated retrieval and acquisition of information.
(2) Knowledge Graph (KG): Compared to search engine, KG is particularly useful for extracting structured, explicit knowledge. Relevant knowledge from a knowledge graph can be extracted and used as a prompt input to enhance the generative process (Sun et al., 2023c). StructGPT (Jiang et al., 2023f) introduced an iterative reading-and-reasoning framework where the model can access a knowledge graph through a well-designed interface, continually acquiring information and reasoning until an answer is obtained. RoG (Luo et al., 2024) generates plausible reasoning paths based on a KG before executing each path in parallel and integrating the outcomes for a final answer. The concept behind ToG (Sun et al., 2023c) is similar to RoG, but instead of pre-planning reasoning paths, it allows the model to explore potential entities and links first and then reason, continuously assessing the feasibility of the reasoning paths.
(3) API-based Tools: An important part of the tools is the real-world APIs, which enable the model to obtain information from specific data sources, such as real-time stock information, movie services, code interpreters and so on. However, the multitude and diversity of APIs, coupled with the adherence to certain operational protocols, make the teaching of API usage to models a focal point of this area. Toolformer (Schick et al., 2023) trains language models in self-supervised manner to automatically call APIs when needed. It commences by utilizing prompts to generate API calls within the text, subsequently executes these calls, and filters out ineffective ones based on the execution results to form the final dataset. By employing standard language modeling objectives on this dataset, it is possible to train a model capable of autonomously invoking APIs across various downstream tasks without losing its inherent language modeling capabilities.
RestGPT (Song et al., 2023) has formulated a comprehensive framework for prompting LLMs to invoke RESTful APIs, comprising an online planner, an API selector, and an executor. ToolLLM (Qin et al., 2023b) capitalizes on a substantial corpus of scraped APIs to build a dataset for fine-tuning. Moreover, Gorilla (Patil et al., 2023) introduces an information retriever to provide the model with reference API documentation, which facilitates the teaching of retrieval-based information utilization during fine-tuning. ToolkenGPT (Hao et al., 2023) incorporates each tool as a new token into the vocabulary. During the training process, the model learns the representation of each token, enabling it to invoke APIs in an inference phase as naturally as generating regular text. Beyond learning to invoke APIs, CREATOR (Qian et al., 2023a) proposes a framework to prompt models to write code based on actual problems as a new tool implementation. The generated tools function through a code interpreter and have demonstrated impressive outcomes on complex mathematical reasoning tasks.
A part of the work additionally supports multimodal inputs, further broadening the application scope of the models. AssistGPT (Gao et al., 2023b) offers a comprehensive framework that includes modules such as Planner, Executor, Inspector, and Learner, utilizing both language and code to enable more intricate inference processes. ViperGPT (Surís et al., 2023), by feeding CodeX with user queries and visual API information, generates corresponding Python code to invoke APIs, successfully completing complex visual tasks.
(4) Model-based Tools: With the swift expansion of diverse AI communities(i.e., Huggingface, ModelScope, Github), various types of AI models have become readily accessible for use, serving as a pivotal tool in enhancing generative retrieval systems. These AI models encompass a wide array of tasks each accompanied by comprehensive model descriptions and usage examples. Owing to their extensive training on specific tasks, these models often exhibit exceptional performance in those tasks. HuggingGPT (Shen et al., 2023) employs ChatGPT as a controller to deconstruct user queries into a sequence of tasks, subsequently determining which models to invoke for task execution. Similarly, Visual ChatGPT (Wu et al., 2023b) integrates a visual foundation model with Large Language Models (LLMs), leveraging ChatGPT as a prompt manager to mobilize various visual foundation models, like BLIP (Li et al., 2022) and ControlNet (Zhang et al., 2023c). It is adept at processing users’ image-based requests and is more efficient compared to multi-modal models.
Figure 6. Generating response with citation and personal information assistant are also crucial approaches for building a reliable and user-centric GenIR system.
4.3. Generating Response with Citation
To build a reliable GenIR system, generating responses with citations is a promising approach (Metzler et al., 2021; Huang and Chang, 2023; Litschko et al., 2023). Citations allow users to clearly understand the source of each piece of knowledge in the response, not only enhancing trust in the GenIR system but also facilitating its widespread adoption. Existing methods can be divided into directly generating responses with citations and using a retrieval module to enhance the generated content.
4.3.1. Direct Generating Response with Citation
This method uses the model’s intrinsic memory to generate source citations without relying on the retrieval module.
(1) Model Intrinsic Knowledge. Leveraging the capabilities of the language model itself, according-to prompting (Weller et al., 2023) guides LLMs to more accurately cite information from the pre-training data to reduce the generation of false information. By adding phrases like ”according to Wikipedia” in the model’s prompts, the model is guided to cite the corresponding knowledge sources.
Furthermore, to improve the quality of citations, several methods are proposed. Iterative Feedback Learning (IFL) (Lee et al., 2023d) first uses a critique model to assess the generated text, then provides targeted feedback based on the assessment results, guiding LLMs to iteratively improve their performance. Through this method, the IFL approach can effectively enhance the accuracy of citations, content correctness, and linguistic fluency of LLMs, while maintaining a high standard of information accuracy. In addition, Fierro et al. (Fierro et al., 2024) proposed a plan-based approach, defining the generation plan as a series of questions that serve as a blueprint for generating content and its organization. They introduced two attribution models utilizing the blueprint: an abstract model, where questions are generated from scratch; and an extractive model, where questions are directly copied from the input. Experiments show that planning consistently improves the quality of citations.
(2) Incorporating Generative Retrieval. As envisioned by Metzler et al. (Metzler et al., 2021), to build an expert-level reliable IR system, allowing the model to directly generate responses with citations is a promising approach. Users do not need to search for answers from a list of returned documents like traditional IR systems but can directly receive reliable responses tailored to their information needs. Moreover, the cited document is also generated by the model through the generative retrieval approach described in Section 3, directly generating corresponding DocIDs.
Utilizing generative retrieval, 1-PAGER (Jain et al., 2023) combines answer generation and evidence retrieval. 1-PAGER gradually generating N-gram DocIDs through constrained decoding using FM-Index (Ferragina and Manzini, 2000), thereby partitioning the retrieval corpus, selecting documents, and generating response step by step. 1-PAGER is comparable to existing retrieval-then-read methods in retrieval and answer accuracy and superior to pure closed-book QA models because it attributes predictions to specific evidence corpora. It provides a new scheme for integrating retrieval into seq2seq response generation.
Recently, (Khalifa et al., 2024) proposes a source-aware training method. Specifically, they teach the model to associate DocIDs with knowledge during pre-training, then provide citations of supporting evidence during instruction tuning. Experiments have shown that this method can effectively achieve knowledge attribution of pre-training data, enhancing the verifiability of LLMs, and providing a promising research direction for building a trustworthy GenIR system.
4.3.2. Retrieval-based Response with Citation
To enhance the accuracy of citations, several methods have been developed based on retrieval techniques to fetch relevant documents, thereby improving the quality of responses with embedded citations.
(1) Citation within Generation. Following retrieval, models directly generate responses that include citations. Initially, systems like WebGPT (Nakano et al., 2021), LaMDA (Thoppilan et al., 2022), and WebBrain (Qian et al., 2023b) were developed. These methods utilize web pages or Wikipedia to construct large-scale pre-training datasets, teaching models how to generate responses with citations.
Subsequently, more advanced strategies for citation generation were proposed. For instance, Search-in-the-Chain (SearChain) (Xu et al., 2023a) introduces a method where it first generates a reasoning chain (Chain-of-Query, CoQ) through prompts to LLMs. The retrieval module interacts with each node of the CoQ for verification and completion tasks. After the interaction, SearChain performs tracing operations, generating the entire reasoning process and marking citations for each step of the inference.
Following retrieval and answer generation, LLatrieval (Li et al., 2023j) suggests continuously improving the retrieval results through an iterative updating process, verifying whether the retrieved documents can adequately support the generated answers until the verification is satisfied. Later, AGREE (Ye et al., 2023) utilizes a Natural Language Inference (NLI) model to verify the consistency between the LLM-generated answers and the retrieved documents, employing a Test-Time Adaptation (TTA) strategy that allows LLMs to actively search and cite the most current information during the generation process, thus enhancing the accuracy and reliability of the responses. VTG (Sun et al., 2023a) integrates an evolved memory system and a dual-layer validator for generating verifiable text, combining long-term and short-term memories to adapt to dynamically changing content focuses and using an NLI model to evaluate the logical support between claims and potential evidences.
Based on the graph of thoughts (GoT) (Besta et al., 2024a), HGOT (Fang et al., 2024) improves context learning in retrieval-augmented settings by constructing a hierarchical GoT. This method leverages the LLM’s planning capabilities to break down complex queries into smaller sub-queries and introduces a scoring mechanism to assess the quality of retrieved paragraphs.
Employing reinforcement learning, Huang et al. (Huang et al., 2024) introduce a fine-grained reward mechanism to train language models, allocating specific rewards for each generated sentence and citation, thus teaching models how to accurately cite external information sources. This approach uses rejection sampling and reinforcement learning algorithms to enhance model performance in generating citation-inclusive text through localized and specialized reward signals. APO (Li et al., 2024d) reimagines the task of attributive text generation as a preference learning problem, automatically generating a large number of preference data pairs to reduce the cost of manual annotation. Through progressive preference optimization and experience replay techniques, this method reinforces the model’s preference signals at a fine-grained level, while avoiding overfitting and text degradation caused by automatic data generation.
(2) Citation after Generation. This approach involves models first generating a response, then adding citations through models like NLI. RARR (Gao et al., 2023a) improves attributability by automatically finding external evidence for the language model’s output and post-editing to correct content while preserving the original output as much as possible. This method combines a few training examples, LLM, and retrieval to enhance its attribution capabilities without altering the existing model. Following that, PURR (Chen et al., 2023c) employs an unsupervised learning method by allowing LLMs to generate text noise themselves, then training an editor to eliminate this noise, achieving an efficient editing process. PURR leverages the generative capabilities of language models to create training data, which not only improves attribution performance but also significantly speeds up the generation process. CEG (Li et al., 2024b) searches for supporting documents related to generated content and uses an NLI-based citation generation module to ensure each statement is supported by citations. ”Attribute First, then Generate” (Slobodkin et al., 2024) decomposes the generation process, first selecting relevant source text details and then generating based on these details, achieving localized attributability in text generation. This method offers a finer granularity of attribution, ensuring high-quality text with each sentence supported by a clear source, greatly reducing the workload of manual fact-checking.
4.4. Personal Information Assistant
The core of the GenIR system is the user, so understanding user intent is crucial. Researchers have explored various methods like personalized search (Zhou et al., 2020; Wang et al., 2023b; Zhou et al., 2024c; liu2023how_to_personalize), dialogue (Zhang et al., 2018; Ma et al., 2021; Liu et al., 2020) and recommender (Liu et al., 2023c; Dai et al., 2023; Zhiyuli et al., 2023) systems to explore the users’ interests. Specifically, personalized information assistants aims to better understand users’ personalities and preferences, generating personalized responses to better meet their information needs. This section reviews the progress in research on personalized dialogue and domain-specific personalization.
4.4.1. Personalized Dialogue System
To better understand user needs, researchers have explored two main approaches: personalized prompt design and model fine-tuning.
(1) Personalized Prompt. For personalized prompt design, Liu et al. (Liu et al., 2023c) and Dai et al. (Dai et al., 2023) input users’ interaction history and rating history into ChatGPT (OpenAI, 2022) for in-context learning, effectively generating personalized responses. LaMP (Salemi et al., 2023) enhances the language model’s personalized output by retrieving personalized history from user profiles. Using long-term history, (Christakopoulou et al., 2023) designs prompts describing users’ long-term interests, needs, and goals for input into LLMs. BookGPT (Zhiyuli et al., 2023) uses LLM prompts, interactive querying methods, and result verification frameworks to obtain personalized book recommendations. PerSE (Wang et al., 2023j) infers preferences from several reviews by a specific reviewer and provides personalized evaluations for new story inputs.
Using prompt rewriting, (Li et al., 2023f) proposes a method combining supervised and reinforcement learning to better generate responses from frozen LLMs. Similarly, (Chen et al., 2023f) rewrites user input prompts using extensive user text-to-image interaction history to align better with expected visual outputs.
(2) Personalized Fine-tuning. This line of work focuses on fine-tuning models for personalized response generation. To generate more engaging dialogues based on users’ personalities and interests, Zhang et al. (Zhang et al., 2018) introduced a dataset with 5 million personas for dialogues, Persona-Chat, to train models to produce more personalized and appealing conversations. Mazaré et al. (Mazaré et al., 2018) created a dataset of over 700 million conversations extracted from Reddit, demonstrating the effectiveness of training dialogue models on a large-scale personal profile dataset. $\mathcal{P}^{2}$ot (Liu et al., 2020) generates more personalized and consistent dialogues by simulating the perception of personalities between conversation participants. DHAP (Ma et al., 2021) designs a novel Transformer model structure to automatically learn implicit user profiles from users’ dialogue history without explicit personal information. Wu et al. (Wu et al., 2021) propose a generative segmentation memory network to integrate diverse personal information. Fu et al. (Fu et al., 2022) developed a variational approach to model the relationship between personal memory and knowledge selection, with a bidirectional learning mechanism allowing mutual learning between personal memory fragments and knowledge selection.
Using reinforcement learning, Cheng et al. (Cheng et al., 2023b) collected a domain-specific preference (DSP) dataset and proposed a three-stage reward model learning scheme, including base language model training, general preference fine-tuning, and customized preference fine-tuning. Jang et al. (Jang et al., 2023a) developed a method called ”Personalized Soups,” first optimizing multiple policy models with different preferences independently using PPO (Schulman et al., 2017), then dynamically combining these models’ parameters during the inference stage.
Using retrieval-enhanced methods, LAPDOG (Huang et al., 2023b) retrieves relevant information from story documents to enhance personal profiles and generate better personalized responses. To effectively utilize multiple knowledge sources, SAFARI (Wang et al., 2023c) leverages LLMs’ capabilities in planning, understanding, and integrating knowledge under supervised and unsupervised training settings, effectively generating responses consistent with character settings and knowledge-enhanced. Inspired by writing education, Li et al. (Li et al., 2023g) proposed a multi-stage, multi-task framework to teach LLMs to generate personalized responses, including retrieval, ranking, summarization, synthesis, and generation. For subjective tasks, (Wozniak et al., 2024) studied the superior performance of personalized fine-tuning in subjective text perception tasks compared to non-personalized models.
To achieve a personalized information assistant for every user, OPPU (Tan et al., 2024c) uses personalized PEFT (Ding et al., 2023) to store user-specific behavioral patterns and preferences, showing superior performance in handling changes in user behavior, modeling users with different activity levels, and among different PEFT methods.
For multimodal scenarios, PMG (Shen et al., 2024) proposes a personalized multi-modal generation method that transforms user behavior into natural language, allowing LLMs to understand and extract user preferences.
4.4.2. Domain-specific Personalization
Understanding users’ personalized information needs, the GenIR system has broad applications across various domains such as healthcare, academia, education, recipes, etc.
(1) Healthcare. In AI-assisted healthcare, personalization plays a crucial role. Liu et al. (Liu et al., 2023d) utilize few-shot tuning to process and infer based on time-series physiological and behavioral data. Zhang et al. (Zhang et al., 2023d) implement specific medical diagnosis identification in databases and prospective diagnostic assistance using prompts from ChatGPT (OpenAI, 2022) and GPT-4 (Achiam et al., 2023). Subsequently, Yang et al. (Besta et al., 2024b) propose an LLM for traditional Chinese medicine called Zhongjing, based on LLaMA (Touvron et al., 2023a), which undergoes a complete training process from continued pre-training, supervised fine-tuning to reinforcement learning with human feedback (RLHF) (Christiano et al., 2017). Abbasian et al. (Abbasian et al., 2023) introduce an open-source LLM-based conversational health agent framework called openCHA, which collects necessary information through specific actions and generates personalized responses. MedAgents (Tang et al., 2023c) propose a multidisciplinary collaboration (MC) framework where LLM-based agents engage in multi-round cooperative discussions through role-playing to enhance the model’s expertise and reasoning capabilities.
For mental healthcare, Mental-LLM (Xu et al., 2023c) presents a new framework for using LLMs to predict mental health from social media text data, with prompting-based and finetuning-based methods for real-time monitoring and prediction of psychological issues such as depression and anxiety. Lai et al. (Lai et al., 2023) introduce a psychological consultation aid called Psy-LLM, combining pre-trained LLMs with real psychologist Q&As and a large corpus of psychological articles.
For medication suggestions, Liu et al. (Liu et al., 2023e) propose a framework called PharmacyGPT for generating personalized patient groups, formulating medication plans, and predicting patient outcomes.
(2) Academic. In the academic domain, RevGAN (Li and Tuzhilin, 2019) can automatically generate controllable and personalized user reviews based on users’ emotional tendencies and stylistic information. For writing assistants, Porsdam et al. (Porsdam Mann et al., 2023) explore the personalized enhancement of academic writing using LLMs like GPT-3 (Brown et al., 2020), showing higher quality in format, style, overall quality, and novelty after training with published academic works of three authors. Similarly, to address the lack of personalized outputs for author communication styles and expertise in current LLM outputs, Mysore et al. (Mysore et al., 2023) propose Pearl, a personalized LLM writing assistant trained on data selected from users’ historical documents to optimize personalized text generation; they also develop a new KL divergence training objective to help retrievers more accurately track document contributions to personalized generation.
(3) Education. In the education domain, Cui et al. (Cui and Sachan, 2023) propose a new adaptive and personalized exercise generation method that dynamically adjusts exercise difficulty to match students’ learning progress by combining knowledge tracing and controlled text generation. EduChat (Dan et al., 2023) learns education-specific functionalities, such as article evaluation and emotional support, through pre-training on educational corpora and fine-tuning on customized instructions, addressing the issues of delayed knowledge updates and lack of educational expertise traditionally faced by LLMs in the educational sector.
(4) Other Domains. For recipe generation tasks, traditional methods fail to consider users’ personal tastes and unfamiliar dishes. To address this, Majumder et al. (Majumder et al., 2019) propose a personalized generation model based on users’ historical recipe consumption. Using an encoder-decoder structure, the model utilizes users’ past recipe data to generate responses, enhancing the personalization and adaptability of the outputs. Subsequently, for personalized headline generation, Zhang et al. (Zhang et al., 2022) simulate users’ interests based on their news browsing history and generate news headlines based on these interests. Salemi et al. (Salemi et al., 2023) propose the LaMP benchmark, which includes various personalized generation tasks such as personalized news headline generation, personalized academic title generation, personalized email subject generation, and personalized tweet rewriting. Additionally, for personalized assistance with home cleaning robots in determining the correct storage locations for items, TidyBot (Wu et al., 2023a) proposes using LLMs to generalize from a small number of user-provided examples to infer broadly applicable user preference rules.
5. Evaluation
This section will provide a range of evaluation metrics and benchmarks for generative information retrieval methods, along with analysis and discussions on their performance.
5.1. Evaluation for Generative Document Retrieval
5.1.1. Metrics
In this section, we will discuss several core metrics for evaluating GR methods, including Recall (Zhu, 2004), R-Precision (Zhu, 2004), Mean Reciprocal Rank (MRR) (Craswell, 2009), Mean Average Precision (MAP) (Zhu, 2004), and Normalized Discounted Cumulative Gain (nDCG) (Järvelin and Kekäläinen, 2002). These metrics provide different perspectives on the effectiveness of a GR system, including its accuracy, efficiency, and the relevance of its results.
Recall (Zhu, 2004) is a metric that measures the proportion of relevant documents retrieved by the search system. For a given cutoff point $k$, the recall $\text{Recall}@k$ is defined as:
(8) $\text{Recall}@k=\frac{1}{|Q|}\sum_{q=1}^{|Q|}\frac{ret_{q,k}}{rel_{q}},$
where $|Q|$ is the number of queries in the set, $ret_{q,k}$ is the number of relevant documents retrieved for the $q$-th query within the top $k$ results, and $rel_{q}$ is the total number of relevant documents for the $q$-th query.
R-Precision (Zhu, 2004) measures the precision at the rank position $R$, which corresponds to the number of relevant documents for a given query $q$. It is calculated as:
[…]
where $ret_{q,R}$ is the number of relevant documents retrieved within the top $R$ positions, and $R$ is equivalent to $rel_{q}$.
MRR (Craswell, 2009) reflects the average rank position of the first relevant document returned in the search results. It is computed as follows:
(10) $\text{MRR}=\frac{1}{|Q|}\sum_{q=1}^{|Q|}\frac{1}{\text{rank}_{q}},$
where $\text{rank}_{q}$ is the rank of the first relevant document returned for the $q$-th query.
MAP (Zhu, 2004) calculates the average precision across multiple queries. It considers the exact position of all relevant documents and is calculated using the following formula:
(11) $\text{MAP}=\frac{1}{|Q|}\sum_{q=1}^{|Q|}\left(\frac{1}{rel_{q}}\sum_{k=1}^{n_{q}}\text{P}@k\times I(q,k)\right),$
where $\text{P}@k$ is the precision at cutoff $k$, and $I(q,k)$ is an indicator function that is 1 if the document at position $k$ is relevant to the $q$-th query and 0 otherwise.
nDCG (Järvelin and Kekäläinen, 2002) takes into account not only the relevance of the documents returned but also their positions in the result list, which is defined by:
(12) $\text{DCG}@k=\sum_{i=1}^{k}\frac{2^{\text{rel}_{i}}-1}{\log_{2}(i+1)},$
(13) $\text{nDCG}@k=\frac{\text{DCG}@k}{\text{IDCG}@k},$
where $\text{rel}_{i}$ represents the graded relevance of the $i$-th document, $\text{DCG}@k$ is the discounted cumulative gain, and $\text{IDCG}@k$ represents the maximum possible $\text{DCG}@k$.
5.1.2. Benchmarks
Evaluating the effectiveness of GR methods relies on high-quality and challenging benchmark datasets. Here are several benchmark datasets that are widely used in the field:
MS MARCO (Nguyen et al., 2016) (Microsoft Machine Reading Comprehension) is a large dataset for evaluating machine reading comprehension, retrieval, and question-answering capabilities in web search scenarios, containing two benchmarks: document ranking and passage ranking, with a total of 3.2 million documents and 8.8 million passages. It is compiled from real user queries extracted from Microsoft Bing’s search logs, each accompanied by annotated relevant documents. This dataset covers a diverse range of question types and document genres, aiming to assess the performance of GR systems in complex web search scenarios.
NQ (Kwiatkowski et al., 2019) (Natural Questions) is a question-answering dataset introduced by Google, using Wikipedia as the corpus, which includes 3.2 million documents, each document being a Wikipedia page. It contains numerous natural user queries and their corresponding answers extracted from web pages in Google search results, which could be utilzed to evaluate the retrieval performance of GR systems in addressing real-world questions.
TriviaQA (Joshi et al., 2017) is a QA dataset that includes a large number of factual questions and answers from various sources, including Wikipedia and web-based QA forums, totaling about 0.66M evidence documents. TriviaQA aims to assess the generalization ability of information retrieval systems across different domains and question types.
KILT (Petroni et al., 2021) (Knowledge Intensive Language Tasks) is a comprehensive benchmark dataset integrating 5 categories of knowledge-intensive tasks, including fact checking (FEVER (Thorne et al., 2018)), entity linking (AIDA CoNLL-YAGO (Hoffart et al., 2011), WNED-WIKI (Guo and Barbosa, 2018), WNED-CWEB (Guo and Barbosa, 2018)), slot filling (T-REx (Elsahar et al., 2018), Zero Shot RE (Levy et al., 2017)), open-domain QA (Natural Questions (Kwiatkowski et al., 2019), HotpotQA (Yang et al., 2018), TriviaQA (Joshi et al., 2017), ELI5 (Fan et al., 2019)), and dialogue (Wizard of Wikipedia (Dinan et al., 2018)). The KILT benchmark also uses Wikipedia as its corpus, comprising 5.9 million wiki pages, which aims to evaluate the effectiveness of information retrieval systems in handling complex language tasks that require extensive background knowledge.
TREC Deep Learning Track 2019 & 2020 (Craswell et al., 2020, 2021) focus on using deep learning to enhance the efficiency of information retrieval, with primary tasks including document and passage ranking. These evaluation campaigns use the MS MARCO dataset to simulate real-world search queries, providing a standardized environment for assessing different retrieval techniques.
It is noteworthy that, due to the challenge of indexing large-scale document corpus by GR models, some methods like (Tay et al., 2022; Zhou et al., 2023b; Wang et al., 2022a; Zhuang et al., 2022; Zhou et al., 2022; Chen et al., 2023a; Jin et al., 2023) have adopted the approach of evaluating retrieval performance on smaller subsets of documents, ranging from 10K to 300K, along with their corresponding training and test pairs.
DynamicIR. For dynamic corpora, DynamicIR (Yoon et al., 2023) proposes a task framework based on StreamingQA (Liska et al., 2022) benchmark for evaluating IR models within dynamically updated corpora. Through experimental analysis, DynamicIR revealed that GR systems are superior in adapting to evolving knowledge, handling temporally informed data, and are more efficient in terms of memory, indexing time, and FLOPs compared to dense retrieval systems.
ExcluIR. For exclusionary retrieval tasks, where users explicitly indicate in their queries that they do not want certain information, ExcluIR (Zhang et al., 2024c) provides a set of resources. This includes an evaluation benchmark and a training set to help retrieval models understand and process exclusionary queries.
5.1.3. Analysis
In addition to the benchmarks and metrics for evaluating the performance of GR methods, there is a series of works that have conducted detailed analyses and discussions to study the behavior of GR models.
Understanding Generative Retrieval. To understand the performance of DSI (Tay et al., 2022) in text retrieval, Chen et al. (Chen et al., 2023a) examines uniqueness, completeness, and relevance ordering. These respectively reflect the system’s ability to distinguish between different documents, retrieve all relevant documents, and accurately rank documents by relevance. Experimental analysis find that DSI excels in remembering the mapping from pseudo queries to DocIDs, indicating a strong capability to recall specific DocIDs from particular queries. However, the study also pointed out DSI’s deficiency in distinguishing relevant documents from random ones, negatively impacting its retrieval effectiveness.
Exploring the connection between generative and dense retrieval, (Nguyen and Yates, 2023) demonstrates that they can be considered as bi-encoders in dense retrieval. Specifically, the authors analyze the computation of dot products during the generative retrieval process, which is similar to the calculation of dot products between query vectors and document vectors in dense retrieval. Following this, (Wu et al., 2024b) revisits generative retrieval from the perspective of multi-vector dense retrieval (MVDR), revealing a common framework in computing document-query relevance between the two methods. This work also analyzes their differences in document encoding and alignment strategies, further confirming through experiments the phenomenon of term matching in the alignment matrices and their commonalities in retrieval.
Large-scale Experimental Analysis. Later, Pradeep et al. (Pradeep et al., 2023) conduct the first comprehensive experimental study on GR techniques over large document sets, such as the 8.8M MS MARCO passages, which results in significant computational demands. It was found that among all the techniques examined, using generated pseudo queries to augment training data remains the only effective method on large document corpus. The strongest result in the experiments was achieved by using a training task that only utilized synthetic queries to Naive DocIDs, expanding the model to T5-XL (3B parameters) to achieve an MRR@10 of 26.7. Surprisingly, increasing the parameters to T5 XXL (11B) in the same setup did not improve performance but rather led to a decline. These findings suggest that more research and in-depth analysis are needed in the GR field, and possibly additional improvements to the paradigm, to fully leverage larger language models.
Out-of-distribution Perspective. For out-of-distribution (OOD) robustness of GR models, Liu et al. (Liu et al., 2023g) investigate three aspects: query variations, new query types, and new tasks. Their study showed that all types of retrieval models suffer from performance drops with query variations, indicating sensitivity to query quality and structure. However, when dealing with new query types and tasks, GR models showed different levels of adaptability, with pre-training enhancing their flexibility. The research highlights the critical need for OOD robustness in GR models for dealing with ever-changing real-world information sources.
5.2. Evaluation for Response Generation
5.2.1. Metrics
Evaluating the quality of generated responses involves multiple aspects, including accuracy, fluency, relevance, etc. In this section, we’ll introduce the main metrics for evaluating reliable response generation, categorized into rule-based, model-based, and human evaluation metrics.
(1) Rule-based Metrics. Exact Match (EM) is a straightforward evaluation method requiring the model’s output to be completely identical to the reference answer at the word level. This full character-level matching is stringent, often used in tasks requiring precise and concise answers, such as question answering systems, e.g., NQ (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), SQuAD (Rajpurkar et al., 2018), etc. It simply calculates the ratio of perfectly matched instances to the total number of instances.
For the generation of longer text sequences, BLEU (Papineni et al., 2002) is a common metric initially used to evaluate the quality of machine translation. It compares the similarity between the model’s output and a set of reference texts by calculating the overlap of n-grams, thereby deriving a score. This method assumes that high-quality generation should have a high lexical overlap with the labeled answer. Optimized from BLEU, METEOR (Banerjee and Lavie, 2005) is an alignment-based metric that considers not only exact word matches but also synonyms and stem matches. Additionally, METEOR introduces considerations for word order and syntactic structure to better assess the fluency and consistency of the generated text.
ROUGE (Lin, 2004), also a commonly used metric for evaluating longer texts, by measuring the extent of overlap in words, sentences, n-grams, and so forth, between the generated text and a collection of reference texts. It focuses on recall, meaning it evaluates how much of the information in the reference text is covered by the generated text. ROUGE comes in various forms, including ROUGE-N, which evaluates based on n-gram overlap, and ROUGE-L, which considers the longest common subsequence, catering to diverse evaluation requirements.
Perplexity (PPL) is a metric for evaluating the performance of language models, defined as the exponentiation of the average negative log-likelihood, reflecting the model’s average predictive ability for a given corpus of text sequences. The lower the perplexity, the stronger the model’s predictive ability. Specifically, given a sequence of words $W=w_{1},w_{2},\ldots,w_{N}$, where $N$ is the total number of words in the sequence, PPL can be expressed as:
(14) $\text{PPL}(W)=\exp\left\{-\frac{1}{N}\sum_{i=1}^{N}\log p(w_{i}|w_{<i})\right\},$
where $p(w_{i}|w_{<i})$ represents the pre-trained language model’s probability of predicting the $i$-th word $w_{i}$ given the previous words $w_{<i}$.
(2) Model-based Metrics. With the rise of pre-trained language models, a series of model-based evaluation metrics have emerged. These metrics utilize neural models to capture the deep semantic relationships between texts.
Unlike traditional rule-based metrics, BERTScore (Zhang et al., 2019) utilizes the contextual embeddings of BERT (Kenton and Toutanova, 2019) to capture the deep semantics of words, evaluating the similarity between candidate and reference sentences through the cosine similarity of embeddings. BERTScore employs a greedy matching strategy to optimize word-level matching and uses optional inverse document frequency weighting to emphasize important words, ultimately providing a comprehensive evaluation through a combination of recall, precision, and F1 score. BERTScore captures not only surface lexical overlap but also a deeper understanding of the semantic content of sentences.
Similarly based on BERT (Kenton and Toutanova, 2019), BLEURT (Sellam et al., 2020) designed multiple pre-training tasks, enhancing the model’s ability to recognize textual differences with millions of synthetic training pairs. These pre-training tasks include automatic evaluation metrics (such as BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), and BERTScore (Zhang et al., 2019)), back-translation likelihood, textual entailment, etc. Each task provides different signals to help the model learn how to evaluate the quality of text generation.
BARTScore (Yuan et al., 2021), based on the pre-trained seq2seq generative model BART (Lewis et al., 2020a), treats the evaluation of generated text as a text generation problem. Specifically, BARTScore determines the quality of text based on the transition probability between the generated text and reference text. BARTScore does not require additional parameters or labeled data and can flexibly evaluate generated text from multiple perspectives (such as informativeness, fluency, factuality, etc.) and further enhance evaluation performance through text prompts or fine-tuning for specific tasks.
FActScore (Min et al., 2023) focuses on the factual accuracy of each independent information point in long texts. It calculates a score representing factual accuracy by decomposing the text into atomic facts and verifying whether these facts are supported by reliable knowledge sources. This method provides a more detailed evaluation than traditional binary judgments and can be implemented efficiently and accurately through human evaluation and automated models (combining retrieval and powerful language models).
GPTScore (Fu et al., 2023) is a flexible, multi-faceted evaluation tool that allows users to evaluate text using natural language instructions without the need for complex training processes or costly annotations. GPTScore constructs an evaluation protocol dynamically through task specification and aspect definition and utilizes the zero-shot capability of pre-trained language models to evaluate text quality, optionally using demonstration samples to improve evaluation accuracy.
(3) Human Evaluation Metrics. Human evaluation is an important method for assessing the performance of language models, especially in complex tasks where automated evaluation tools struggle to provide accurate assessments. Compared to rule-based and model-based metrics, human evaluation is more accurate and reliable in real-world applications. This evaluation method requires human evaluators (such as experts, researchers, or everyday users) to provide comprehensive assessments of the model-generated content based on their intuition and knowledge.
Human evaluation measures the quality of language model outputs by integrating multiple assessment criteria, following (Chang et al., 2023): Accuracy (Singhal et al., 2022) primarily evaluates the correctness of information and its correspondence with facts; Relevance (Zhong et al., 2022) focuses on whether the model’s output is pertinent to the specific context and user query; Fluency (van der Lee et al., 2019) examines whether the text is coherent, natural, and facilitates smooth communication with users; Safety (Ji et al., 2023b) scrutinizes whether the content may lead to potential adverse consequences or harm. These indicators collectively provide a comprehensive assessment of the model’s performance in real-world settings, ensuring its effectiveness and applicability.
5.2.2. Benchmarks and Analysis
In this section, we explore various benchmarks for evaluating the performance of language models in generating reliable responses. These benchmarks assess language understanding, factual accuracy, reliability, and the ability to provide timely information.
(1) General Evaluation. To comprehensively assess the language models’ understanding capabilities across a wide range of scenarios, MMLU (Hendrycks et al., 2021) utilizes a multiple-choice format covering 57 different tasks, from basic mathematics to American history, computer science, and law. This benchmark spans evaluations in humanities, social science, and science, technology, engineering, and mathematics, providing a comprehensive and challenging test. It has been widely used in the evaluation of Large Language Models (LLMs) in recent years (Touvron et al., 2023a, b; Jiang et al., 2023b).
Furthermore, BIG-bench (Srivastava et al., 2022) introduces a large-scale and diverse benchmark designed to measure and understand the capabilities and limitations of LLMs across a broad range of tasks. Including 204 tasks contributed by 450 authors from 132 institutions, it covers areas such as linguistics, mathematics, and common sense reasoning. It focuses on tasks beyond the capabilities of language models, exploring how model performance and societal biases evolve with scale and complexity.
LLM-Eval (Lin and Chen, 2023) offers a unified multi-dimensional automatic evaluation method for open-domain dialogue of LLMs, eliminating the need for manual annotation. The performance of LLM-Eval across various datasets demonstrates its effectiveness, efficiency, and adaptability, improving over existing evaluation methods. The research also analyzes the impact of different LLMs and decoding strategies on the evaluation outcomes, underscoring the importance of selecting suitable LLMs and decoding strategies.
For Chinese, C-Eval (Huang et al., 2023a) aims to comprehensively evaluate LLMs’ advanced knowledge and reasoning capabilities in the Chinese context. It is based on a multiple-choice format, covering four difficulty levels and 52 different academic fields from secondary school to professional levels. C-Eval also introduces C-Eval Hard, a subset containing highly challenging subjects to test the models’ advanced reasoning capabilities. Through evaluating state-of-the-art English and Chinese LLMs, C-Eval reveals areas where current models still fall short in handling complex tasks, guiding the development and optimization of Chinese LLMs.
(2) Tool Evaluation. To assess the ability of language models to utilize tools, API-Bank (Li et al., 2023i) provides a comprehensive evaluation framework containing 73 APIs and 314 tool usage dialogs, along with a rich training dataset of 1,888 dialogs covering 1,000 domains to improve LLMs’ tool usage capabilities. Experiments show that different LLMs perform variably in tool usage, highlighting their strengths and areas for improvement.
Later, ToolBench (Qin et al., 2023b) developed a comprehensive framework including a dataset and evaluation tools to facilitate and assess the ability of LLMs to use over 16,000 real-world APIs. It enhances reasoning capabilities by automatically generating diverse instruction and API usage scenario paths, introducing a decision tree based on depth-first search. ToolBench significantly enhances LLMs’ performance in executing complex instructions and in their ability to generalize to unseen APIs. ToolLLaMA, an LLM fine-tuned from LLaMA (Touvron et al., 2023a), exhibits remarkable zero-shot capabilities and performance comparable to state-of-the-art LLMs like ChatGPT (OpenAI, 2022).
(3) Factuality Evaluation. TruthfulQA (Lin et al., 2022) measures the truthfulness of language models in answering questions. This benchmark consists of 817 questions covering 38 categories, including health, law, finance, and politics. This evaluation reveals that, even in optimal conditions, the truthfulness of model responses only reaches 58%, in stark contrast to human performance at 94%. Moreover, they proposed an automated evaluation metric named GPT-judge, which classifies the truthfulness of answers by fine-tuning the GPT-3 (Brown et al., 2020) model, achieving 90-96% accuracy in predicting human evaluations.
HaluEval (Li et al., 2023a) is a benchmark for evaluating LLM illusions, constructed using a dataset containing 35K illusion samples, employing a combination of automated generation and manual annotation. This provides effective tools and methods for assessing and enhancing large language models’ capabilities in identifying and reducing illusions. For Chinese scenarios, HalluQA (Cheng et al., 2023a) designs 450 meticulously selected adversarial questions to assess the illusion phenomenon in Chinese LLMs, covering multiple domains and reflecting Chinese culture and history, identifying two main types of illusions: imitative falsehoods and factual errors.
To evaluate the ability of LLMs to generate answers with cited text, ALCE (Gao et al., 2023c) builds an end-to-end system for retrieving relevant text passages and generating answers with citations. ALCE contains three datasets, covering different types of questions, and evaluates the generated text’s quality from ’fluency’, ’correctness’, and ’citation quality’ dimensions, combining human evaluation to verify the effectiveness of the evaluation metrics. The experimental results show that while LLMs excel at generating fluent text, there is significant room for improvement in ensuring content factual correctness and citation quality, especially on the ELI5 dataset where the best model was incomplete in citation support half of the time.
(4) Real-Time Evaluation. RealTime QA (Kasai et al., 2023) created a dynamic question-and-answer platform that regularly releases questions and evaluates systems weekly to ask and answer questions about the latest events or information. It challenges the static assumption of traditional QA datasets aiming for immediate application. Experiments based on LLMs like GPT-3 and T5 found that models could effectively update their generated results based on newly retrieved documents. However, when the retrieved documents failed to provide sufficient information, models tended to return outdated answers.
Furthermore, FreshQA (Vu et al., 2023) evaluates large language models’ performance in challenges involving time-sensitive and erroneous premise questions by creating a new benchmark containing questions of this nature. Evaluating various open and closed-source LLMs revealed significant limitations in handling questions involving rapidly changing knowledge and erroneous premises. Based on these findings, the study proposed a simple in-context learning method, FreshPrompt, significantly improving LLMs’ performance on FreshQA by integrating relevant and up-to-date information sourced from search engines into the prompt.
(5) Safety, Ethic, and Trustworthiness. To comprehensively evaluate the safety of LLMs, SafetyBench (Zhang et al., 2023a) implements an efficient and accurate evaluation of LLMs’ safety through 11,435 multiple-choice questions covering 7 safety categories in multiple languages (Chinese and English). The diversity of question types and the broad data sources ensure rigorous testing of LLMs in various safety-related scenarios. Comparing the performance of 25 popular LLMs, SafetyBench revealed GPT-4’s significant advantage and pointed out the areas where current models need improvements in safety to promote the rapid development of safer LLMs.
For ethics, TrustGPT (Huang et al., 2023e) aims to assess LLMs’ ethical performance from toxicity, bias, and value alignment, three key dimensions. The benchmark uses predefined prompt templates based on social norms to guide LLMs in generating content and employs multiple metrics to quantitatively assess the toxicity, bias, and value consistency of these contents. Experimental analysis revealed that even the most advanced LLMs still have significant issues and potential risks in these ethical considerations.
For trustworthiness, TrustLLM (Sun et al., 2024) explores principles and benchmarks including truthfulness, safety, fairness, robustness, privacy, and machine ethics across six dimensions. Extensive experiments, including assessing 16 mainstream LLMs’ performance on 30 datasets, found that trustworthiness usually positively correlates with functional effectiveness. While proprietary models typically outperform open-source models in trustworthiness, some open-source models like Llama2 showed comparable high performance.
These benchmarks provide important tools and metrics for evaluating and improving the capabilities of language models, contributing to the development of more accurate, reliable, safe, and timely GenIR systems. For further understanding of the evaluation works, (Wang et al., 2023f; Chang et al., 2023; Huang et al., 2023d; Dai et al., 2024) offer more detailed introductions.
6. Challenges and Prospects
This section discusses the key challenges faced in the fields of generative document retrieval and reliable response generation, as well as potential directions for future research.
6.1. Challenges on Generative Document Retrieval
6.1.1. Scalability Issues
As extensively studied by (Pradeep et al., 2023), generative retrieval demonstrates significantly lower retrieval accuracy compared to dense retrieval when handling million-level document corpora in web search scenarios. Merely increasing the model size does not yield stable performance improvements. However, GR outperforms dense retrieval in document collections smaller than 300K, posing a question: What impedes GR methods from scaling to large document sizes? This issue encompasses several aspects:
Training Data. Current LLMs are pre-trained on huge datasets ranging from hundreds of billions to several trillion tokens, covering vast knowledge sources such as the internet, books, and news articles, consuming substantial computational power (Zhao et al., 2023). They are then extensively fine-tuned with high-quality, human-annotated data to achieve substantial generalization capabilities (Raffel et al., 2020; Lewis et al., 2020a; Ouyang et al., 2022; Touvron et al., 2023a). In contrast, generative retrieval (GR) models often begin with a pre-trained language model and are fine-tuned on labeled data comprising ¡query, DocID¿ pairs, which does not sufficiently prepare them to fully grasp GR tasks. For numeric-based DocIDs, the models, having not encountered these numbers in their pre-training phase, tend to rote memorize the DocIDs seen during training, struggling to predict unseen ones effectively. Similarly, if text-based DocIDs fail to precisely represent the documents, the model also tend to rote learning.
A potential solution is to create a large-scale pre-training dataset for generative retrieval on a general corpus, possibly including a variety of common DocIDs such as URLs, titles, and numerical sequences. We can utilize instructions to distinguish generation targets for various DocIDs. Then we can pre-train a Transformer-based GR model from scratch, the model can understand generative retrieval across diverse domains. This method could bridge the gap between language model pre-training data and GR tasks, enhancing the generalization ability of GR models across different corpora.
Training Method. As described in Section 3.1.1, existing training methods explore various training objectives, including seq2seq training, learning DocID, and ranking capabilities. Other methods involve knowledge distillation (Chen et al., 2023b), reinforcement learning (Zhou et al., 2023a), etc. Is there a better training method to enable GR models to master generating DocID ranking lists? For example, RLHF (Christiano et al., 2017) has been effectively used to train LLMs (Ouyang et al., 2022; Touvron et al., 2023b), though at a high cost. Exploring RLHF in the GR field is also worthwhile.
Model Structure. As discussed in Section 3.1.2, most current GR models are based on encoder-decoder Transformers structures (Tay et al., 2022; Wang et al., 2022a; Zhou et al., 2022), such as T5 (Raffel et al., 2020) and BART (Lewis et al., 2020a). Some GR methods like CorpusLM (Li et al., 2024a) have experimented with a decoder-only structure of the LLM Llama2 (Touvron et al., 2023b), requiring more training computational power but not significantly improving performance. Research is needed to determine which structure is more suitable for generative retrieval. Additionally, whether increasing model and data size could lead to emergent phenomena similar to those observed in LLMs (Wei et al., 2022a; Schaeffer et al., 2023) is also a promising research direction.
6.1.2. Handling Dynamic Corpora
Real-world applications often involve dynamically changing corpora, such as the web and news archives, where incremental learning is essential. However, for language models, indexing new documents inevitably leads to forgetting old ones, posing a challenge for GR systems. Existing methods like DSI++ (Mehta et al., 2022), IncDSI (Kishore et al., 2023), CLEVER (Chen et al., 2023d), and CorpusBrain++ (Guo et al., 2024) propose solutions such as experience replay, constrained optimization, incremental product quantization, and continual generative pre-training frameworks to address incremental learning issues. Yet, these methods have their specific applicable scenarios, and more effective and universally applicable incremental learning strategies remain a key area for exploration.
6.1.3. Document Identifier
Accurately representing a document with high-quality DocIDs is crucial for generative retrieval.
For example, the KILT dataset based on the Wikipedia corpus, which includes 5.9 million documents, demonstrates optimistic retrieval performance for GR methods using titles as DocIDs (Cao et al., 2021b; Chen et al., 2022b; Li et al., 2024a). This is because each document in Wikipedia has a unique manually annotated title that represents the core entity discussed in that page. However, in the web search scenario, such as in the MS MARCO dataset (Nguyen et al., 2016), many documents lack a unique title, are overlapping, and the titles do not accurately represent the core content of the documents. Thus, GR performance significantly declines in the MS MARCO corpus of 8.8 million passages.
Therefore, how to construct high-quality titles (or other types of DocIDs) in general corpora, similar to those in Wikipedia, that not only accurately represent documents but also are lightweight, is a critical factor for implementing GR methods and warrants in-depth research.
Text or Numeric? As discussed in Section 3.2, current methods include text-based and numeric-based DocIDs, each with their advantages and disadvantages. Text-based DocIDs effectively leverage the linguistic capabilities of pre-trained generative language models and offer better interpretability. Numeric-based DocIDs can utilize dense retriever embeddings to obtain semantic DocID sequences; they can also complement dense retrievers to achieve synergistic benefits.
However, to ensure good generalization ability of GR models without extensive pre-training, it is essential to utilize the inherent pre-trained parameters of the model. Coherent textual DocIDs can naturally leverage this aspect, but they also need to capture key document semantics and maintain linguistic sequence characteristics. Numeric DocIDs, however, do not offer this advantage. Thus, as mentioned in 6.1.1, extensive pre-training is necessary to enable models to fully understand the meanings behind these numerical strings, which is a costly endeavor.
Do We Need a Unique ID for Each Document? Most current GR methods use a unique DocID to uniquely identify a document. However, as the number of documents in a corpus increases, maintaining a unique DocID becomes increasingly challenging. Even if a unique DocID is maintained, it is difficult to differentiate significantly from other DocIDs semantically, leading to reduced retrieval precision. Some methods, such as using sub-string as DocIDs (Bevilacqua et al., 2022; Chen et al., 2023e), have proven effective. These methods utilize the FM-Index (Ferragina and Manzini, 2000) to ensure the generated sub-string exists in the corpus and use the number of generated sub-strings in different documents to rank documents, demonstrating good performance and generalization ability.
However, since this method is based on FM-Index, its inference latency is high, which is an issue that needs addressing. Furthermore, exploring other more efficient alternatives to FM-Index and even considering not using constrained search but freely generating a DocID sequence followed by a lightweight matching and scoring module to efficiently return a document ranking list are also worthy of exploration.
6.1.4. Efficiency Concerns
Current GR methods generally rely on constrained beam search to generate multiple DocID sequences during inference, resulting in high latency. This is particularly severe when returning 100 or more documents, with latencies reaching several hundred milliseconds (Wang et al., 2022a), which is unacceptable for low-latency IR systems. Therefore, designing more efficient inference methods is crucial. To reduce inference latency, the length of the DocID sequence should not be too long; 16 tokens or fewer is an efficient range. This necessitates designing DocIDs that are precise and concise enough to represent documents while maintaining performance and improving efficiency. Additionally, developing more efficient decoding strategies is a valuable research direction for the future.
6.2. Challenges on Reliable Response Generation
6.2.1. Improving Accuracy and Factuality
In GenIR systems, ensuring content accuracy and factuality is crucial. To achieve this, as mentioned in Section 4, there are two main areas of improvement:
Internal Knowledge Memorization. Firstly, training stronger generative models is critical for building reliable GenIR systems. Various commercial LLMs continue to progress, utilizing vast training data and computational resources, but exploring better model structures is also worthwhile. Recent research such as Retentive Networks (Sun et al., 2023b), Mamba (Gu and Dao, 2023), and others, have shown potential to challenge the performance and efficiency of Transformers (Vaswani et al., 2017). However, whether these can scale and truly surpass Transformer-based LLMs in generation quality is still an open question. Moreover, what types of training data and methods can consistently produce models capable of generating high-quality, reliable text also deserve thorough investigation and summary. The mechanisms by which language models recall knowledge during inference are not yet clear and need to be fully understood to better serve user information needs.
External Knowledge Enhancement. As described in Section 4.2.1, retrieval-augmented generation is an effective method widely applied in LLMs. However, there is still room for improvement. For example, whether inserting retrieved documents directly into generative models via prompts is the best method, or if there are better ways, such as inputting embeddings (Yen et al., 2024), needs exploration. Additionally, whether models can autonomously decide whether to perform retrieval (Wang et al., 2024c; Tan et al., 2024a), and when in the generation process to perform it (Jiang et al., 2023e), are topics worth further exploration.
Tool-augmented generation, as discussed in Section 4.2.2, is also a popular method for endowing LLMs with fine-grained world knowledge and performing complex tasks. Recent research has raised questions, such as ”Should tools always be used?” (Wang et al., 2024a). More specifically, whether the performance improvements brought by using tools justify the extra computational costs incurred during model training or the inference costs during testing. Existing work mainly focuses on task accuracy, but studying the cost-effectiveness of these methods is also a valuable topic.
6.2.2. Real-time Properties of GenIR Systems
Timeliness is critical for GenIR systems, as well as traditional IR systems, to provide users with the most up-to-date information. However, since the knowledge of pre-trained generative models is fixed after training, methods like retrieval and tool augmentation are needed to acquire new external knowledge. Research on real-time knowledge acquisition remains limited, making it a valuable area for investigation.
Moreover, continually relying on outdated knowledge from language models is inadequate, as models cannot comprehend the significance of given contexts or backgrounds in the current era, thus reducing the reliability of the generated content. Therefore, updating the information in language models while avoiding the forgetting of existing knowledge, such as through continual learning (Wu et al., 2022; Wang et al., 2023k), knowledge editing (Wang et al., 2023m; Ovadia et al., 2023b; Yao et al., 2023a; Mazzia et al., 2023), etc., is a topic worth further exploring.
6.2.3. Bias and Fairness
Since LLMs are often trained on large, unfiltered datasets, GenIR systems may propagate stereotypes and biases present in the data regarding race, culture, and other aspects (Gallegos et al., 2023). Researchers have explored various methods to enhance the fairness of generated content during training data selection, training methods, generation techniques, and rewriting phases. However, biases have not been eradicated and require a thorough understanding of the mechanisms by which generative models produce biases, to design methods to solve them and build fair GenIR systems that further the practical application of GenIR.
6.2.4. Privacy and Security
Firstly, the content generated by GenIR systems risks plagiarism (Dien, 2023; Kenthapadi et al., 2023). For instance, studies such as (Carlini et al., 2021; Huang et al., 2022) indicate that pre-trained language models can reproduce large segments of their training data, leading to inadvertent plagiarism and causing academic dishonesty or copyright issues. On one hand, legal regulations regarding the copyright of AI-generated content will gradually emerge and evolve. On the other hand, technical research aimed at reducing plagiarism by generative models, such as generating text with correct citations (Metzler et al., 2021; Liu et al., 2023f; Huang and Chang, 2023), is a promising research direction for reliable GenIR that has received increasing attention in recent years.
Moreover, due to the unclear mechanisms of memory and generation in pre-trained language models, GenIR systems inevitably return unsafe content. For example, (Carlini et al., 2019, 2021) show that when attacked, LLMs may return private information of users seen in training data. Therefore, understanding the mechanisms by which LLMs recall training data and designing effective defense mechanisms to enhance security are crucial for the widespread use of GenIR systems. Additionally, developing effective detection methods for content generated by LLMs is essential for enhancing the security of GenIR systems (Yang et al., 2023b).
6.3. Unified Framework
This article discusses two mainstream forms of GenIR: generative document retrieval and reliable response generation. However, each approach has its advantages and limitations. Generative document retrieval still returns a list of documents, whereas the reliable response generation model itself cannot effectively capture document-level relationships. Therefore, integrating these two approaches is a promising research direction.
6.3.1. Unified Framework for Retrieval and Generation
Given that both generative retrieval and downstream generation tasks can be based on generative language models, could a single model perform both retrieval and generation tasks? Indeed, it could.
Current attempts, such as UniGen (Li et al., 2024g), use a shared encoder and two decoders for GR and QA tasks respectively, and show superior performance on small-scale retrieval and QA datasets. However, they struggle to generalize across multiple downstream tasks and to integrate with powerful LLMs. Additionally, CorpusLM (Li et al., 2024a) uses a multi-task training approach to obtain a universal model for GR, QA, and RAG. Yet, merely merging training data does not significantly improve retrieval and generation performance, and CorpusLM remains limited to the Wikipedia corpus. Facing a broader internet corpus presents significant challenges.
In the future, can we construct a large search model (LSM) that allows an LLM to have the capability to generate DocIDs and reliable responses autonomously? Even LSM could decide when to generate DocIDs to access the required knowledge before continuing generation. Unlike the large search model defined in (Wang et al., 2023i), which unifies models beyond the first-stage retrieval (such as re-ranking, snippet, and answer models), we aim to integrate the first-stage retrieval as well, enabling the LSM to fully understand the meaning of retrieval and its connection with various downstream generation tasks.
6.3.2. Towards End-to-End Framework for Various IR Tasks
As envisioned by Metzler et al. (Metzler et al., 2021), propose an expert-level corpus model that not only possesses linguistic capabilities but also understands document-level DocIDs and knows the sources of its own knowledge. Such a model could not only solve the issue of hallucinations common in traditional language models but could also generate texts with references pointing to the source documents, thus achieving a reliable end-to-end GenIR model. By understanding DocIDs and knowledge sources, this end-to-end system could also perform additional IR tasks, such as returning the main content of a document given its DocID or returning other related document DocIDs, as well as enabling multi-lingual and multi-modal retrieval.
Current methods, as discussed in this GenIR survey, primarily focus on generative document retrieval (GR) and response generation as separate entities. GR models excel at comprehending document identifiers at the document-level, while downstream models demonstrate powerful task generation capabilities. However, existing methods face challenges when it comes to effectively integrating these two generative abilities, limiting the overall performance and effectiveness of the GenIR system. The integration of these generative abilities in a seamless and efficient manner remains a key challenge in the field.
In the future, we can design training methods that align knowledge and DocIDs and construct high-quality training datasets for generating answers with references, to train such an end-to-end GenIR model. Achieving this goal remains challenging and requires the collaborative efforts of researchers to contribute to building the next generation of GenIR systems.
7. Conclusion
In this survey, we explore the latest research developments, evaluations, current challenges, and future directions in generative information retrieval (GenIR). We discuss two main directions in the GenIR field: generative document retrieval (GR) and reliable response generation. Specifically, we systematically review the progress of GR covering model training, document identifier design, incremental learning, adaptability to downstream tasks, multi-modal GR, and generative recommendation systems; as well as advancements in reliable response generation in terms of internal knowledge memorization, external knowledge enhancement, generating responses with citations, and personal information assistance. Additionally, we have sorted out the existing evaluation methods and benchmarks for GR and response generation. We organize the current limitations and future directions of GR systems, addressing scalability, handling dynamic corpora, document representation, and efficiency challenges. Furthermore, we identify challenges in reliable response generation, such as accuracy, real-time capabilities, bias and fairness, privacy, and security. We propose potential solutions and future research directions to tackle these challenges. Finally, we also envision a unified framework, including unified retrieval and generation tasks, and even building an end-to-end framework capable of handling various information retrieval tasks. Through this review, we hope to provide a comprehensive reference for researchers in the GenIR field to further promote the development of this area.